### v1

# Batch speculative decoding Done right

Abstract
Speculative decoding speeds up LLM inference by using a small draft model to propose multiple tokens that a target model verifies in parallel. Extending this idea to batches is essential for production serving, but it introduces the ragged tensor problem: sequences in the same batch accept different numbers of draft tokens, breaking right-alignment and corrupting position IDs, attention masks, and KV-cache state. We show that several existing batch implementations violate output equivalence—the fundamental requirement that speculative decoding must produce identical token sequences to standard autoregressive generation. These violations occur precisely due to improper handling of the ragged tensor problem. In response, we (1) characterize the synchronization requirements that guarantee correctness, (2) present a correctness-first batch speculative decoding EqSpec that exposes realignment as consuming 40% of overhead, and (3) introduce EXSpec, which maintains a sliding pool of sequences and dynamically forms same-length groups, to reduce the realignment overhead while preserving per-sequence speculative speedups. On the SpecBench dataset, across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B target/draft pairs, our approach achieves up to 3× throughput improvement at batch size 8 compared to batch size 1, with efficient scaling through batch size 8, while maintaining 95% output equivalence. Our method requires no custom kernels and integrates cleanly with existing inference stacks. Our code is available at https://github.com/eBay/spec_dec.
1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) accelerates LLM inference by using a small draft model to propose multiple tokens that the target model verifies in parallel, shifting from memory-bound sequential generation to compute-intensive verification and delivering single-sequence speedups (Xia et al., 2024). Batch speculative decoding aims to combine this per-sequence acceleration with standard batching by processing multiple sequences (batch dimension) while verifying multiple draft tokens per sequence (sequence dimension). The core challenge is the ragged-tensor effect (Qian et al., 2024): in each verification round, sequences accept a single series of the proposed draft tokens, but prefix lengths differ across sequences (e.g., one accepts five tokens while another accepts one), misaligning sequence lengths. This raggedness violates the rectangular-tensor assumption required for GPU-parallel execution.
Figure 1 highlights a critical, often overlooked issue in batch speculative decoding: methods with impressive throughput can produce corrupted outputs. For lossless acceleration, speculative decoding must yield identical outputs to standard autoregressive generation (Leviathan et al., 2023). As a reference point, the widely used HuggingFace implementation (Wolf et al., 2020) preserves this guarantee, but only for batch size 1. By contrast, in our tests of public batch implementations—specifically, BSP (Su et al., 2023) and DSD (Yan et al., 2025)—this requirement is violated at batch sizes >1>1, manifesting as repetitive tokens or <unk> symbols under greedy decoding rather than matching standard generation.
YES, High Throughput…Refer to caption
BUT, Output Gibberish!
Method Generated Output
Prompt: ”The weather today is”
Ground Truth not as bad as it was yesterday…
Spec-1 ✓ Matches ground truth
BSP ✗ not not not not not not not…
DSD ✗ ⟨unk⟩⟨unk⟩ The perfect time…
EqSpec ✓ Matches ground truth
EXSpec ✓ Matches ground truth
Figure 1: Batch speculative decoding on Vicuna-7B/68M: Existing methods achieve high throughput but violate the fundamental requirement of output equivalence by producing corrupted outputs. Our approach maintains perfect correctness while still achieving competitive performance.
These failures share a single root cause—broken synchronization invariants (position tracking, attention, KV-cache) across ragged tensors. We first propose EqSpec by formalizing these invariants and enforcing them with synchronization-aware scheduling (Section 3.1). Concretely, EqSpec specifies the minimal synchronization needed for correctness and shows that realignment accounts for 40% of computation—an inherent cost of maintaining invariants across ragged tensors. This fundamental overhead helps explain why even production systems struggle: vLLM’s speculative decoding underperforms its non-speculative baseline at higher batch sizes (leading to deprecation in its v1 engine), while SGLang with EAGLE (Li et al., 2024a) consistently exhibits negative speedups. Our analysis indicates that the superlinear growth of synchronization overhead is an inevitable cost of correctness in batch speculative decoding—a barrier no existing system has overcome.
To address this in practice, our main algorithm EXSpec (Section 3.2) expands the scheduling scope: it maintains a sliding window of active sequences and dynamically groups those with identical lengths, eliminating realignment for homogeneous groups. This strategy preserves the scaling efficiency of standard batching—where GPU parallelism drives throughput—while retaining the per-sequence acceleration benefits of speculation. At batch size 8, we achieve a $3\times$ throughput improvement over batch size 1. Nevertheless, beyond batch size 8, throughput degrades as grouping success rates decline, forcing more frequent fallbacks to expensive realignment. Section 4 examines these scaling dynamics and the relationship between sequence diversity, grouping effectiveness, and alignment overhead.
Our experiments on SpecBench (Xia et al., 2024) yield two main results. First, Correctness: unlike prior approaches such as BSP and DSD, which suffer severe output corruption, our method preserves approximately 95% output equivalence across Vicuna-7B/68M (Zheng et al., 2023), Qwen3-8B/0.6B (Yang et al., 2025), and GLM-4-9B/0.6B (GLM et al., 2024) model pairs. Second, Scalability: at batch size 8, EXSpec achieves up to a $3\times$ speedup over batch size 1. Our contributions are summarized as follows:
  • We provide a correctness-first analysis of the ragged-tensor problem in batch speculative decoding, identifying precise synchronization requirements for correctness and explaining why existing methods fail (Section 2).
  • We present a unified solution that maintains correctness through precise synchronization invariants while avoiding their overhead via cross-batch scheduling of same-length sequence groups (Section 3).
  • We experimentally demonstrate that our approach achieves both >>95% output correctness and positive scaling through batch size 8, successfully multiplying batch parallelism with per-sequence speculation gains, whereas production systems (vLLM, SGLang) exhibit negative speedups (Section 4).
2 Design Space Analysis
Figure 2: The ragged tensor problem in batch speculative decoding. Differing numbers of accepted draft tokens across sequences in the same batch lead to ragged-shaped input IDs tensors and KV Cache that disrupt subsequent batch operations.
When sequences within a batch accept different numbers of draft tokens during verification, tensors become irregular, violating GPUs’ requirement for rectangular layouts—this is the ragged-tensor problem illustrated in Figure 2. Despite batching’s centrality to production deployments, existing implementations lack a principled design that preserves output equivalence with standard decoding while scaling with batch size. We identify three approaches to handle raggedness: Masking, Rollback, and Dynamic Padding. Yet, as our systematic analysis shows, current instantiations of these approaches fail to simultaneously maintain correctness and performance at scale. To close this gap, we first analyze the pitfalls of each approach and then introduce a correctness-first algorithmic design with explicit synchronization requirements for reliable batch speculative decoding.
✗Masking Approach (non-contiguous position IDs).
This approach operates directly on ragged tensors by masking rejected tokens in attention and reassigning position IDs so new tokens align with their content positions. Across verification rounds with varying rejections, sequences accumulate padding in various positions (middle and right), forming non-contiguous position IDs that standard Transformer implementations handle poorly. BSP (Su et al., 2023) attempts this via masking but fails to maintain position-ID consistency across iterations, yielding corrupted outputs (Figure 1). EAGLE’s experimental batching code [^1] While EAGLE’s main contribution concerns improved draft models rather than batching, its repository includes experimental batch-related code we analyzed for implementation challenges. https://github.com/SafeAILab/EAGLE/issues/250 (Li et al., 2025) encounters similar framework limitations. Supporting non-contiguous position IDs would require custom CUDA kernels for each base model (Qian et al., 2024)—a prohibitive engineering cost that sacrifices portability.
✗ Rollback Approach (speculation waste).
After each verification step, all sequences are truncated to the batch’s minimum accepted length (Wolf et al., 2020; kamilakesbi, 2024). This guarantees alignment but discards correctly verified tokens from faster sequences. As batch size grows and acceptance-rate variance widens, the waste compounds; in the extreme, one persistently rejecting sequence forces single-token progress for the entire batch. In effect, throughput collapses to that of the worst-performing sequence, undermining speculative gains and rendering the approach impractical at larger scales.
Algorithm 1 BatchVerify: Single Forward Pass
1: Target model $\mathcal{M}_{t}$, Sequences $\mathcal{S}$, Draft tokens $D$, KV cache
2: Accepted tokens AA, Bonus tokens $B$
3: if first iteration then
4:   $\mathcal{X}\leftarrow\mathcal{S}\oplus D$
5: else
6:   $\mathcal{X}\leftarrow D$
7: $\textit{logits},\textit{KVCache}\leftarrow\mathcal{M}_{t}(\mathcal{X},\textit{KVCache})$
8: $\textit{pred\_tokens}\leftarrow\arg\max(\textit{logits, dim=vocab})$
9: $\triangleright$ Vectorized first mismatch detection
10: $\textit{matches}\leftarrow(\textit{pred\_tokens}=D)$
11: $J\leftarrow\arg\max(\neg\textit{matches, dim=seq})$
12: $\triangleright$ Ragged shape acceptance, no vectorization
13: for each sequence $i$ in batch do
14:   $A[i]\leftarrow D[i][:j]$
15:   $\triangleright$ Get bonus token from first mismatch
16:   $\textit{bonus\_logit}\leftarrow\textit{logits}[i,|\mathcal{S}[i]|+j]$
17:   $B[i]\leftarrow\arg\max(\textit{bonus\_logit})$
18: return $A,B,KVCache$
Algorithm 2 EqSpec
1: Draft model $\mathcal{M}_{d}$, Target model $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Max tokens $T$, Draft length $K$
2: Generated sequences $\mathcal{S}$
3: $\mathcal{S}\leftarrow\texttt{Tokenize}(\mathcal{P})$ $\triangleright$ Batch left padding
4: $\textit{KVCache}\leftarrow\emptyset$
5: while until max new tokens do
6:   Phase 1: Draft Generation
7:   $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{S},K)$
8:   Phase 2: Batch Verification
9:   $A,B,\textit{KVCache}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{S},D,\textit{KVCache})$
10:   $\triangleright$ See Figure 3 for illustration on index offset.
11:   Phase 3: Unpad-Append-Repad
12:   for each sequence $i$ in batch do
13:    $\mathcal{S}[i]\leftarrow\texttt{Unpad}(\mathcal{S}[i])$
14:    $\mathcal{S}[i]\leftarrow\mathcal{S}[i]\oplus A[i]\oplus B[i]$
15:   $\mathcal{S},\textit{offset}\leftarrow\texttt{BatchRepad}(\mathcal{S})$
16:   $\textit{KVCache}\leftarrow\texttt{Realign}(\textit{KVCache},\textit{offset})$
17: return $\mathcal{S}$
✓ Dynamic Padding Approach.
This approach realigns sequences after each verification by adjusting left padding to maintain right alignment, preserving all accepted tokens. While conceptually simple, correctness requires tight synchronization of position IDs, attention masks, and the KV-cache. DSD’s experimental code (Yan et al., 2025) follows this idea but merely repads at each step—adding varying left padding without ever unpadding—thereby inflating sequences. It also contains three critical errors: (i) sampling bonus tokens from the draft model rather than the target model; (ii) redundantly regenerating KV-cache entries, causing memory bloat; and (iii) desynchronizing padding, position IDs, and the KV-cache across iterations. Despite the overhead of repeated realignment, a correct dynamic-padding implementation fits within standard frameworks and preserves all verified tokens.
Among the three approaches, only dynamic padding is viable: position-ID schemes require custom kernels that undermine portability, and rollback discards verified tokens at rates that grow with batch size; dynamic padding maintains correctness within standard frameworks.
3 Method
We present a synchronization-aware approach to batch speculative decoding that co-designs correctness and efficiency. The core tension is that preserving correctness requires synchronizing position IDs, attention masks, and the KV-cache across ragged tensors—an overhead that can consume 40% of computation. We introduce two complementary mechanisms: EqSpec specifies and enforces the minimal synchronization invariants required for valid computation (Section 3.1), while EXSpec groups same-length sequences to avoid synchronization overhead (Section 3.2). Together, these form a unified system in which correctness constraints drive scheduling, enabling both output equivalence and practical performance.
Figure 3: EqSpec synchronizes via unpad–append–repad.
3.1 Minimal Batch Realignment: EqSpec
Figure 3 illustrates the core challenge and our remedy for maintaining correctness in batch speculative decoding. After each verification round, sequences accept different numbers of draft tokens, producing ragged tensors that GPUs cannot process. To restore a valid batch, we apply an unpad–append–repad procedure that converts ragged outputs back to a rectangular layout while preserving three invariants: contiguous position IDs, valid attention masks, and aligned KV-cache entries.
Correct implementation requires padding-agnostic position IDs that reset to zero at the first content token, attention masks that exclude padding tokens, and precise KV-cache realignment after each verification step (Algorithm 1). After unpad–append–repad, index offsets shift across sequences; consequently, position IDs and attention masks must be recomputed to preserve correct token relationships. Moreover, the bonus token introduces a special case: it is sampled from the target model’s output distribution at the first mismatch position after the verification forward pass has already completed. Consequently, while it encodes the target’s authoritative correction, it lacks KV-cache entries in either the draft or target model because KV-cache is computed only for tokens that were present in the input sequence during the forward pass—and the bonus token, being an output, was not yet part of that input sequence. Therefore, it must be included in the next forward pass to create its KV-cache entries, further complicating synchronization. Finally, the realignment is resource-intensive: the KV-cache consists of rank-4 tensors $(\text{batch}\times\text{heads}\times\text{sequence}\times\text{dimension})$, and each padding adjustment triggers allocation and concatenation of high-dimensional zeros. Algorithm 2 details the complete EqSpec procedure.
Figure 4: EXSpec pools ragged sequences by length, avoiding realignment; only unmatched sequences need syncing, turning fixed overhead into optional cost.
Speedup Analysis.
The speedup of speculative decoding depends on both the token-acceptance rate and computational costs. The original formulation by Leviathan et al. (2023) analyzes single-sequence performance, modeling the expected tokens generated per iteration as $(1-\alpha^{k+1})/(1-\alpha)$, where $\alpha$ is the token acceptance rate (TAR) and $k$ is the number of draft tokens per speculation round. This single-sequence view assumes negligible batch overhead and focuses purely on acceptance dynamics.
Batch speculative decoding, however, introduces additional complexities not captured by that model—most notably alignment overhead, which can dominate and degrade performance. We therefore introduce a batch-aware speedup model:
$S\;=\;\frac{\alpha\cdot k}{\,c_{\text{draft}}+c_{\text{verify}}+c_{\text{overhead}}(B)\,}$ (1)
where $S$ denotes speedup relative to non-speculative decoding (i.e., running the target model alone), $c_{\text{draft}}$ and $c_{\text{verify}}$ are the relative costs of draft generation and target verification, and $c_{\text{overhead}}(B)$ captures batch-dependent alignment overhead absent from single-sequence analysis.
Algorithm 3 EXSpec: Cross-Batch Scheduling
1: Draft and Target model $\mathcal{M}_{d}$, $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Window size $W$, Batch size $B$
2: Generated sequences $\mathcal{S}$
3: $\textit{Pool}\leftarrow\texttt{InitSequencePool}(\mathcal{P})$
4: $\triangleright$ Tokenize and optionally sort by length
5: $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
6: while Pool.hasActive() do
7:   Phase 1: Lazy Realignment
8:   $\mathcal{B},\textit{mask},\textit{KV}\leftarrow\texttt{GetBatch}(\textit{Window},B)$
9:   $\triangleright$ Try same-length concatenation
10:   $\triangleright$ Fallback to Unpad-Repad Realignment
11:   Phase 2: Draft Generation
12:   $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{B},\textit{mask},K)$
13:   Phase 3: Batch Verification
14:   $A,B,\textit{KV}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{B},D,\textit{KV})$
15:   Phase 4: Write-Back and Window Refill
16:   for $i\in\mathcal{B}$ do
17:    $\textit{Pool}[i]\leftarrow\textit{Pool}[i]\oplus A[i]\oplus B[i]$
18:    $\textit{Pool.KV}[i]\leftarrow\textit{KV}[i]$
19:    if $\textit{isComplete}(\textit{Pool}[i])$ then $\textit{Pool.deactivate}(i)$
20:   $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
21: return Pool.sequences
Two observations follow. First, when sequences within a batch share the same length, batch realignment overhead is negligible; likewise, with very small draft models or prompt lookahead decoding (Fu et al., 2024b), $c_{\text{draft}}$ becomes small relative to verification. Second, and critically, $c_{\text{overhead}}(B)$ scales superlinearly with batch size $B$ due to the ragged-tensor problem (Section 4.3). This overhead decomposes as $c_{\text{overhead}}(B)=c_{\text{pad}}(B)+c_{\text{kv}}(B)$, where $c_{\text{pad}}(B)$ accounts for unpad–append–repad operations and $c_{\text{kv}}(B)$ for KV-cache realignment on rank-4 tensors. While $c_{\text{draft}}$ and $c_{\text{verify}}$ benefit from GPU parallelism and remain relatively stable within the hardware’s batch-parallel regime, the alignment overhead grows with both batch size and variance in acceptance rates across sequences.
3.2 Cross-Batch Scheduling: EXSpec
Profiling EqSpec shows that alignment overhead consumes 39.4% of computation at batch size 8, rising to 46.7% at batch size 16. Because this cost is inherent to synchronizing ragged tensors, micro-optimizing the primitives yields limited gains. Instead of accelerating these operations, EXSpec avoids them by scheduling.
EXSpec differs from EqSpec in how it manages sequence lifecycles. Rather than maintaining fixed batches that require realignment after every verification, we introduce a SequencePool that holds sequences individually in their ragged states. This enables three optimizations: (i) sequences that complete (reach EOS) are immediately removed, avoiding wasted computation on finished items; (ii) lazy realignment defers synchronization until strictly necessary, keeping sequences in their natural ragged form between steps; and (iii) dynamic batch formation over a sliding window of $W>B$ sequences greatly increases the chance of finding same-length groups that can be concatenated without any realignment.
Figure 4 illustrates the flow. After verification, ragged sequences return directly to the pool without realignment. The scheduler then scans the active window and attempts to form batches of identical length. Same-length sequences concatenate directly—no padding adjustments, no position-ID recomputation, no KV-cache realignment—bypassing $c_{\text{overhead}}(B)$. Only when same-length grouping fails do we fall back to the expensive unpad–append–repad procedure from EqSpec. Algorithm 3 formalizes this in four phases: dynamic batch formation (prioritizing same-length groups), draft generation, verification, and pool write-back with continuous window refresh. Combined with prompt-length sorting, grouping rates approach unity for similar workloads, turning the worst case (constant realignment) into the rare case.
4 Experiments
We evaluate EqSpec and EXSpec along two dimensions often conflated in prior work: correctness (whether outputs match non-speculative generation) and throughput (tokens per second). Through systematic evaluation across three model families and comparisons with both research prototypes and production systems, we show that our approach uniquely preserves output equivalence while achieving competitive speedups.
4.1 Experimental Setup
Models.
To demonstrate generality, we evaluate three target–draft pairs: Vicuna-7B/68M (Zheng et al., 2023), Qwen3-8B/0.6B (Yang et al., 2025), and GLM-4-9B/0.6B (GLM et al., 2024). Unless otherwise noted, experiments use NVIDIA A100 80GB GPUs, PyTorch 2.7, HuggingFace Transformers 4.51.3, five draft tokens per speculation round, and greedy decoding for determinism.
Evaluation and Datasets.
We use SpecBench (Xia et al., 2024), focusing on the first turn of each conversation because our evaluation targets offline batch inference rather than multi-turn interaction. We also use Multi30k (Elliott et al., 2016) for a controlled EXSpec study that contrasts random sampling with an identical-length subset, isolating sequence-length diversity as the driver of grouping rate. For the main evaluation, we measure: (1) Throughput: tokens/s across batch sizes; and (2) Correctness: exact-match rate (full-sequence equivalence with non-speculative decoding) and partial-match rate (fraction of tokens matching until the first divergence). The partial-match metric helps localize failure modes—early divergence typically indicates position-ID or KV-cache misalignment.
Batch Speculative Decoding Compared.
Following our design-space taxonomy (Section 2), we evaluate: (1) Masking approaches: BSP (Su et al., 2023) attempts masking with adaptive speculation but suffers position-ID inconsistencies (BASS (Qian et al., 2024) also follows this approach but requires custom CUDA kernels, limiting generality); (2) Dynamic-padding approaches: DSD (Yan et al., 2025) explores padding but mishandles the KV-cache, while EqSpec implements correct synchronization and EXSpec adds cross-batch scheduling; (3) Reference baselines: Spec-1 (batch-size-1 speculation from Hugging Face Transformers), which does not support batch speculative decoding [^2] https://github.com/huggingface/transformers/issues/32165. We also compare with production systems vLLM [^3] We use the vLLM v0 engine because v1 deprecates speculative decoding. (Kwon et al., 2023) (which subsumes TETRIS (Wu et al., 2025) and TurboSpec (Liu et al., 2025) as vLLM forks) and SGLang-EAGLE [^4] SGLang is compatible only with the EAGLE family; we compare Vicuna-7B/EAGLE2 and Qwen3-8B/EAGLE3. There are no available weights for GLM-4. (Zheng et al., 2024; Li et al., 2024b; 2025), noting that their continuous batching and memory-management designs complicate direct comparison. We further test SGLang-EAGLE-Deterministic (SGLang Team, 2025), which enables deterministic execution to reduce numerical variance. No existing implementation uses rollback due to its inherent wastefulness.
Method Vicuna Qwen3 GLM4
Batch 1 Batch 4 Batch 1 Batch 4 Batch 1 Batch 4
E P E P E P E P E P E P
Batch-based Methods (vs. batch=1 non-spec)
Non-Spec-Batch 53.8 98.2 92.9 96.5 93.3 97.2
Spec-1 97.1 98.4 94.6 97.2 96.0 98.0
EqSpec 97.3 98.6 92.1 98.6 94.6 96.9 92.3 95.7 96.7 98.1 96.5 98.3
EXSpec 97.3 98.6 90.8 97.6 94.6 96.9 95.0 97.1 96.7 98.1 95.2 97.7
DSD 0.0 8.1 0.0 2.2 0.2 2.2 0.0 0.6 0.0 1.0 0.0 0.8
BSP 1.9 39.7 0.2 31.3 3.5 19.9 2.1 12.6 1.0 15.3 0.6 8.1
Continuous Batching Systems (vs. own non-spec)
vLLM + Spec 96.9 / 98.0 65.6 / 78.5 72.7 / 84.5
SGLang + EAGLE 69.8 / 79.5 47.7 / 65.4
SGLang + EAGLE + Det 85.0 / 90.0 50.6 / 69.5
Table 1: Correctness check reports exact match (E) and partial match (P) scores. Top: compares with batch=1 non-spec baseline. Bottom: with each system’s non-speculative mode (scores reported as E / P). Our approach sustains $95\%$ accuracy, while prior work (DSD, BSP) suffers major drops from KV-cache and position ID errors.
Figure 5: Decomposing batch speculative decoding performance. (a) Batch scaling efficiency normalized to BS=1, isolating GPU parallelism from per-sequence speculation; EXSpec initially exceeds the No-Ragged-Scaling upper bound before degrading due to alignment overhead. (b) Alignment overhead grows super-linearly with batch size, consuming up to 38% of inference time, validating that $c_{\text{overhead}}(B)$ dominates at scale. (c) Cross-batch grouping rates on Multi30k for random vs. uniform-length sequences, showing that length homogeneity transforms grouping effectiveness.
4.2 Output Correctness Verification
We verify correctness using deterministic greedy decoding to eliminate sampling variance and enable precise bug isolation. This avoids metrics such as ROUGE (Qian et al., 2024), which can mask implementation failures (e.g., repetitive corruption can still score reasonably). Instead, we use exact match (any divergence) and partial match (fraction of tokens before the first mismatch) to diagnose failure modes.
Table 1 reveals distinct patterns. Our methods maintain $\approx$95% exact match across settings, the remaining $\approx$5% divergence stems from numerical non-determinism in floating-point operations (He & Lab, 2025) and tie-breaking in argmax sampling rather than algorithmic errors. By contrast DSD and BSP fail catastrophically with different signatures. DSD’s near-zero scores indicate immediate position-ID misalignment—the model fails from the first token. BSP shows higher partial match (up to 39.7%) but low exact match, indicating gradual degradation: outputs are initially correct before KV-cache drift misdirects attention, triggering repetition. These complementary patterns—immediate failure vs. gradual decay—show how partial-match metrics reveal not just that implementations fail, but when and why.
Production systems (vLLM, SGLang) introduce additional complexity. While both are accurate on Vicuna, they degrade markedly on Qwen3. SGLang-EAGLE-Deterministic helps disambiguate the cause: improved accuracy suggests most divergences stem from floating-point non-determinism (He & Lab, 2025) rather than algorithmic bugs. Despite this, we show below that their throughput still falls below non-speculative baselines.
4.3 Overhead and Scaling Dynamics
We now validate our theoretical predictions through five studies that decompose batch speculative decoding performance. These experiments isolate alignment overhead growth, quantify its impact on batch scaling, identify sequence diversity as the key bottleneck to grouping effectiveness, compare against production systems, and provide mechanistic insights through overhead profiling.
Batch scaling efficiency.
Figure 1 shows that EXSpec outperforms EqSpec in absolute throughput by combining speculative and batching gains, yet EqSpec exhibits negative scaling beyond BS=8. To test whether batch speculative decoding can retain GPU-parallelism benefits despite raggedness, Figure 5(a) normalizes each method’s throughput to its BS=1 baseline, isolating batch-level scaling from per-sequence speculation and enabling comparison to No-Ragged-Scaling (standard batched decoding without speculation). A notable effect emerges: EXSpec initially surpasses even this upper bound because speculation converts memory-bound operations into compute-bound verification, improving GPU utilization when managed correctly. However, at larger batch sizes, alignment overhead dominates and the advantage inverts, confirming that $c_{\text{overhead}}(B)$ eventually overwhelms parallelism gains.
Figure 6: Speculative decoding lags non-speculative baselines, with larger batches further degrading throughput.
Alignment overhead growth.
[…]
Grouping rate $\times$ sequence-length distribution.
Figure 5(c) probes whether cross-batch scheduling limits are algorithmic or circumstantial via Multi30k. Comparing random sampling to an All-Mean subset with identical lengths isolates the bottleneck: sequence diversity, not the method. Random sampling shows grouping rates collapsing with batch size, whereas the All-Mean configuration maintains high grouping effectiveness even at moderate scales, substantially reducing $c_{\text{overhead}}(B)$. This contrast suggests that preprocessing strategies (e.g., bucketing, dynamic sorting) can push real workloads toward the ideal, revealing untapped potential when scheduling is paired with workload shaping.
Production systems limitation.
Figure 6 shows that production frameworks struggle with batch speculative decoding. Although our method is not directly comparable to continuous-batching systems (which vary effective batch size dynamically), the trend is consistent: vLLM’s speculative decoding underperforms its baseline at high concurrency (leading to deprecation in v1), and SGLang+EAGLE is slower than non-speculative generation across all batch sizes. Two architectural factors likely contribute: (i) speculation repurposes the prefill stage for parallel verification, breaking the prefill–decode separation these systems optimize; and (ii) continuous batching already induces raggedness from varying request lengths, and speculation compounds this with per-sequence acceptance variance, creating nested raggedness that overwhelms existing schedulers. These outcomes indicate that speculation cannot simply be grafted onto architectures optimized for predictable token-by-token decoding; deeper redesign is required.
Method TPS $|\text{Spec}|$ Time/Verif. (ms) Verif. % Draft % Overhead %
EqSpec 95.6 1469 24.88 44.9 27.4 27.7
EXSpec 156.4 952 29.13 55.9 29.5 14.6
Table 2: Overhead anatomy of batch speculation methods. Despite slower per-verification time, EXSpec achieves higher throughput by reducing total verification calls and minimizing alignment overhead through intelligent scheduling.
Overhead Anatomy.
Table 2 shows that EXSpec attains higher throughput via a deliberate trade-off. Cross-batch scheduling cuts total verification calls by one-third and halves alignment overhead by grouping same-length sequences. The trade-off is memory locality: dynamic batching scatters KV-cache entries, increasing per-verification latency. Despite slower individual operations, the reduction in operation count yields a 64% overall throughput gain. This balance between operation count and efficiency suggests future work on improving KV-cache locality within the cross-batch framework.
5 Related Work
Speculative Decoding.
Speculative decoding accelerates LLM inference by verifying draft tokens in parallel. Two verification paradigms dominate: sequence verification—as in SpecDec and follow-ups (Xia et al., 2023; Santilli et al., 2023; Yang et al., 2023; Hooper et al., 2023; Zhang et al., 2023; Fu et al., 2024a)—which preserves the target model’s output distribution; and tree verification—e.g., Medusa/EAGLE variants (Miao et al., 2024; Spector & Re, 2023; Sun et al., 2023; He et al., 2023; Cai et al., 2024; Li et al., 2024a; b; 2025)—which explores multiple branches to raise acceptance rates. Recent works parallelize multiple draft sequences per request (Stewart et al., 2024; Lee et al., 2024) but still verify each request independently. Approximate schemes (Kim et al., 2023; Zhong et al., 2025) trade exactness for speed; our work assumes lossless verification to detect implementation errors rather than approximation artifacts.
Batch Speculative Decoding.
Extending speculative decoding to batch settings introduces the ragged tensor problem: when sequences accept different numbers of draft tokens, the resulting variable-length tensors break GPU-friendly rectangular operations (Qian et al., 2024). Existing approaches face fundamental limitations detailed in Section 2. Position ID/masking approaches like BSP (Su et al., 2023) suffer from position ID inconsistencies across iterations. Dynamic padding approaches reveal critical implementation errors: both DSD (Yan et al., 2025) and Meta’s recent work (Tang et al., 2025) incorrectly sample bonus tokens from the draft model’s distribution rather than the target model’s, violating the fundamental correctness guarantee of speculative decoding. This error compounds with improper KV-cache handling, producing corrupted outputs—BSP generates repetitive tokens while DSD produces <unk> symbols. BASS (Qian et al., 2024) sidesteps these issues through custom CUDA kernels but sacrifices portability. These failures demonstrate that the ragged tensor problem extends beyond performance to correctness itself, motivating our cross-batch scheduling approach that maintains correctness while achieving batch scaling without custom kernels.
Production Systems for Speculative Decoding.
Serving stacks integrate speculation atop general batching (vLLM, SGLang/EAGLE, and forks such as TETRIS and TurboSpec (Kwon et al., 2023; Zheng et al., 2024; Li et al., 2025; Wu et al., 2025; Liu et al., 2025)). In practice, variable draft acceptance clashes with continuous batching and paged attention, and our measurements indicate scaling limits (e.g., small single-sequence gains that reverse at higher concurrency); some engines have since reduced or deprecated speculative-decoding support. We intentionally build our reference implementation from scratch, prioritizing correctness checks; compatibility with continuous batching, paged attention, and prefill–decode separation is orthogonal and left as future work.
6 Conclusion
We presented a correctness-first approach to batch speculative decoding that resolves the fundamental tension between maintaining output equivalence and achieving practical speedups. By identifying the precise synchronization invariants required for correctness, we explained why existing methods fail and established the minimal requirements for valid computation. In particular, our EqSpec implementation enforces these invariants but reveals that realignment overhead consumes up to 40% of computation. By contrast, EXSpec overcomes this limitation via cross-batch scheduling that dynamically groups same-length sequences to bypass realignment entirely. Moreover, experiments across three model families show that our approach uniquely preserves >>95% output equivalence while achieving up to 3× throughput improvement at batch size 8, hereby validating that a careful co-design of correctness constraints and scheduling can multiply batch parallelism with per-sequence speculation gains. Our findings on ragged shape tensor synchronization overhead and cross-batch scheduling may inform future designs for production systems, where integrating speculation with continuous batching remains an open challenge.
Acknowledgment
This work is supported by an eBay research grant.
Appendix A Appendix
Reproducibility Statement
We provide our complete implementation in the Supplementary Material, including all hyperparameters, experimental configurations, and model specifications detailed in Section 4. Our correctness verification framework using exact and partial match metrics enables deterministic validation of both our results and future implementations. All experiments use publicly available models and datasets for reproducibility.
Model Configurations.
We evaluate three target-draft model pairs: Vicuna-7B (lmsys/vicuna-7b-v1.3) with a 68M draft model (double7/vicuna-68m), Qwen3-8B (Qwen/Qwen3-8B) with Qwen3-0.6B (Qwen/Qwen3-0.6B), and GLM-4-9B (zai-org/GLM-4-9B-0414) with a 0.6B draft model (jukofyork/GLM-4.5-DRAFT-0.6B-v3.0). All models were loaded in FP16 precision with greedy decoding (temperature=0, top_p=1.0) to ensure deterministic outputs. We use five draft tokens per speculation round across all experiments unless otherwise specified.
Production Systems Configuration.
We evaluated two production inference systems for comparison: vLLM version 0.9.1 and SGLang commit c4e314f (the deterministic decoding had not merged to the stable version at the time of submission, so we compiled it from source). For vLLM, we used the V0 engine with speculative decoding enabled, as the V1 engine does not support draft model speculative decoding [^5] https://github.com/vllm-project/vllm/issues/21797. For SGLang, we used EAGLE-based speculation with model-specific draft models: yuhuili/EAGLE-Vicuna-7B-v1.3 for Vicuna-7B and Tengyunw/qwen3_8b_eagle3 for Qwen3-8B. GLM-4 was not evaluated with SGLang due to the unavailability of compatible EAGLE draft models. We tested both standard SGLang inference and SGLang with deterministic mode enabled (SGLang Team, 2025) to isolate floating-point non-determinism from algorithmic correctness issues.
Disclose on LLM usage
We used Large Language Models as assistive tools in preparing this manuscript: GPT-5 and Claude Opus 4.1 for polishing writing (grammar correction and sentence restructuring), generating conceptual diagram illustrations, and writing unit tests; NVIDIA/Parakeet-TDT-0.6B-v3 for voice input transcription. LLMs were not used for research ideation, experimental design, data analysis, or scientific conclusions. All core algorithmic implementations and scientific contributions are solely from the authors. We take full responsibility for all content in this paper.
Ethics Statement
This work focuses on improving the efficiency of LLM inference without altering model outputs or introducing biases. Our correctness-preserving approach ensures that acceleration techniques do not compromise model safety or reliability. Our open-source release enables reproducible research and transparent evaluation of batch speculative decoding techniques.

### v2

@@
 Abstract
-Speculative decoding speeds up LLM inference by using a small draft model to propose multiple tokens that a target model verifies in parallel.
-Extending this idea to batches is essential for production serving, but it introduces the ragged tensor problem: sequences in the same batch accept different numbers of draft tokens, breaking right-alignment and corrupting position IDs, attention masks, and KV-cache state.
-We show that several existing batch implementations violate output equivalence—the fundamental requirement that speculative decoding must produce identical token sequences to standard autoregressive generation.
-These violations occur precisely due to improper handling of the ragged tensor problem.
-In response, we (1) characterize the synchronization requirements that guarantee correctness, (2) present a correctness-first batch speculative decoding EqSpec that exposes realignment as consuming 40% of overhead, and (3) introduce EXSpec, which maintains a sliding pool of sequences and dynamically forms same-length groups, to reduce the realignment overhead while preserving per-sequence speculative speedups.
-On the SpecBench dataset, across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B target/draft pairs, our approach achieves up to 3× throughput improvement at batch size 8 compared to batch size 1, with efficient scaling through batch size 8, while maintaining 95% output equivalence.
-Our method requires no custom kernels and integrates cleanly with existing inference stacks.
-Our code is available at https://github.com/eBay/spec_dec.
+Speculative decoding must produce outputs distribution identical to standard autoregressive generation—this output equivalence is not an optimization target but the defining criterion of valid speculative decoding.
+We demonstrate that all existing batch speculative decoding implementations violate this fundamental requirement, producing corrupted outputs ranging from repetitive tokens to gibberish.
+These failures stem from the ragged tensor problem: sequences in the same batch accept different numbers of draft tokens, desynchronizing position IDs, attention masks, and KV-cache state.
+We present the first authentic batch speculative decoding framework.
+We (1) formalize the synchronization invariants that valid batch speculative decoding must satisfy, (2) present EQSPEC, the first algorithm that guarantees output equivalence, and analyze its cost structure to show that alignment overhead grows superlinearly and consumes up to 40% of computation, and (3) introduce EXSPEC, which reduces this overhead through cross-batch scheduling that dynamically groups same-length sequences.
+On SpecBench across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B pairs, our methods achieve up to 3× throughput improvement at batch size 8 while maintaining algorithmic correctness.
+Our methods achieve 95% decoding-equivalence, with residual divergence attributable to floating-point non-determinism in GPU inference, not the synchronization failures that cause near-zero equivalence of prior methods.
+Our code is available at https://github.com/eBay/spec_dec
+Large Language Models, Speculative Decoding, Inference Optimization, Batch Processing
 1 Introduction
-Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) accelerates LLM inference by using a small draft model to propose multiple tokens that the target model verifies in parallel, shifting from memory-bound sequential generation to compute-intensive verification and delivering single-sequence speedups (Xia et al., 2024).
-Batch speculative decoding aims to combine this per-sequence acceleration with standard batching by processing multiple sequences (batch dimension) while verifying multiple draft tokens per sequence (sequence dimension).
-The core challenge is the ragged-tensor effect (Qian et al., 2024): in each verification round, sequences accept a single series of the proposed draft tokens, but prefix lengths differ across sequences (e.g., one accepts five tokens while another accepts one), misaligning sequence lengths.
-This raggedness violates the rectangular-tensor assumption required for GPU-parallel execution.
-Figure 1 highlights a critical, often overlooked issue in batch speculative decoding: methods with impressive throughput can produce corrupted outputs.
-For lossless acceleration, speculative decoding must yield identical outputs to standard autoregressive generation (Leviathan et al., 2023).
-As a reference point, the widely used HuggingFace implementation (Wolf et al., 2020) preserves this guarantee, but only for batch size 1.
-By contrast, in our tests of public batch implementations—specifically, BSP (Su et al., 2023) and DSD (Yan et al., 2025)—this requirement is violated at batch sizes >1>1, manifesting as repetitive tokens or <unk> symbols under greedy decoding rather than matching standard generation.
+Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) accelerates LLM inference by using a small draft model to propose multiple tokens that the target model verifies in parallel.
+The fundamental requirement of speculative decoding is output equivalence (Leviathan et al., 2023; Xia et al., 2024): the generated token distribution must remain identical to standard autoregressive generation.
+Any method that violates this requirement, regardless of its throughput, does not constitute valid speculative decoding.
+Batch speculative decoding aims to combine this per-sequence acceleration with standard batching, but as we demonstrate, all existing implementations fail to preserve output equivalence when batch size exceeds one.
+Figure 1 highlights a fundamental but systematically overlooked failure in batch speculative decoding: methods with impressive throughput can produce corrupted outputs.
+The widely used HuggingFace implementation (Wolf et al., 2020) preserves output equivalence, but only for batch size 1.
+By contrast, in our tests of public batch implementations—specifically, BSP (Su et al., 2023) and DSD (Yan et al., 2025), produce repetitive tokens or <unk> symbols under greedy decoding rather than matching standard generation.
+These are not minor quality degradations but fundamental violations of decoding equivalence.
+Equivalence to reference decoding is a yes/no property: prior batch implementations fail to preserve the target model’s output distribution, yielding near-zero exact-match to standard autoregressive decoding due to synchronization errors.
+The root cause is improper handling of the ragged tensor problem (Qian et al., 2024), which desynchronizes position IDs, attention masks, and KV-cache state across sequences in a batch.
 YES, High Throughput…Refer to caption
@@
 Spec-1 ✓ Matches ground truth
-BSP ✗ not not not not not not not…
-DSD ✗ ⟨unk⟩⟨unk⟩ The perfect time…
+BSP not not not not not not not…
+DSD ⟨unk⟩⟨unk⟩ The perfect time…
 EqSpec ✓ Matches ground truth
@@
 Figure 1: Batch speculative decoding on Vicuna-7B/68M: Existing methods achieve high throughput but violate the fundamental requirement of output equivalence by producing corrupted outputs.
-Our approach maintains perfect correctness while still achieving competitive performance.
-These failures share a single root cause—broken synchronization invariants (position tracking, attention, KV-cache) across ragged tensors.
-We first propose EqSpec by formalizing these invariants and enforcing them with synchronization-aware scheduling (Section 3.1).
-Concretely, EqSpec specifies the minimal synchronization needed for correctness and shows that realignment accounts for 40% of computation—an inherent cost of maintaining invariants across ragged tensors.
-This fundamental overhead helps explain why even production systems struggle: vLLM’s speculative decoding underperforms its non-speculative baseline at higher batch sizes (leading to deprecation in its v1 engine), while SGLang with EAGLE (Li et al., 2024a) consistently exhibits negative speedups.
-Our analysis indicates that the superlinear growth of synchronization overhead is an inevitable cost of correctness in batch speculative decoding—a barrier no existing system has overcome.
-To address this in practice, our main algorithm EXSpec (Section 3.2) expands the scheduling scope: it maintains a sliding window of active sequences and dynamically groups those with identical lengths, eliminating realignment for homogeneous groups.
-This strategy preserves the scaling efficiency of standard batching—where GPU parallelism drives throughput—while retaining the per-sequence acceleration benefits of speculation.
-At batch size 8, we achieve a $3\times$ throughput improvement over batch size 1.
-Nevertheless, beyond batch size 8, throughput degrades as grouping success rates decline, forcing more frequent fallbacks to expensive realignment.
-Section 4 examines these scaling dynamics and the relationship between sequence diversity, grouping effectiveness, and alignment overhead.
+Our approach preserves reference-decoding equivalence while still achieving competitive performance.
+Valid batch speculative decoding requires maintaining synchronization invariants (position IDs, attention masks, KV-cache) across ragged tensors, a challenge that prior implementations fail to address.
+We formalize these invariants, prove they are preserved by our update rule, and present EqSpec, the first algorithm that guarantees output distribution equivalence by enforcing them after each verification round (Section 3.1).
+We analyze the cost structure of batch speculative decoding and show that alignment overhead grows superlinearly with batch size (Section 3.2); empirically, realignment consumes up to 40% of computation at batch size 8 (Section 4.3).
+This is an inherent cost of algorithmic correctness across ragged tensors, not an implementation inefficiency.
+The superlinear growth of alignment overhead represents a fundamental barrier that explains why prior methods either sacrifice validity or fail to scale.
+To reduce this overhead, we introduce EXSpec (Section 3.2), which expands the scheduling scope beyond a single batch.
+EXSpec maintains a pool of active sequences and dynamically groups those with identical lengths, eliminating realignment for homogeneous groups.
+This strategy preserves the scaling efficiency of standard batching while retaining per-sequence speculation benefits.
+At batch size 8, EXSpec achieves a 3× throughput improvement over batch size 1.
+Beyond batch size 8, throughput degrades as sequence length diversity reduces grouping effectiveness, forcing more frequent fallbacks to realignment.
+Section 4.3 examines these scaling dynamics in detail.
 Our experiments on SpecBench (Xia et al., 2024) yield two main results.
-First, Correctness: unlike prior approaches such as BSP and DSD, which suffer severe output corruption, our method preserves approximately 95% output equivalence across Vicuna-7B/68M (Zheng et al., 2023), Qwen3-8B/0.6B (Yang et al., 2025), and GLM-4-9B/0.6B (GLM et al., 2024) model pairs.
-Second, Scalability: at batch size 8, EXSpec achieves up to a $3\times$ speedup over batch size 1.
-Our contributions are summarized as follows:
-  • We provide a correctness-first analysis of the ragged-tensor problem in batch speculative decoding, identifying precise synchronization requirements for correctness and explaining why existing methods fail (Section 2).
-  • We present a unified solution that maintains correctness through precise synchronization invariants while avoiding their overhead via cross-batch scheduling of same-length sequence groups (Section 3).
-  • We experimentally demonstrate that our approach achieves both >>95% output correctness and positive scaling through batch size 8, successfully multiplying batch parallelism with per-sequence speculation gains, whereas production systems (vLLM, SGLang) exhibit negative speedups (Section 4).
+First, Algorithmic Correctness: unlike prior approaches such as BSP and DSD, which suffer severe output corruption, our method maintains algorithmic correctness across Vicuna-7B/68M (Zheng et al., 2023), Qwen3-8B/0.6B (Yang et al., 2025), and GLM-4-9B/0.6B (GLM et al., 2024) model pairs, achieving 95% decoding-equivalence with residual divergence attributable to floating-point non-determinism in GPU inference, not synchronization failures.
+Second, Scalability: at batch size 8, EXSpec achieves up to a $3\times$ speedup over batch size 1, successfully combining batch parallelism with per-sequence speculation gains.
+Our contributions:
+  • We identify that existing batch speculative decoding implementations violate output equivalence due to improper handling of the ragged tensor problem.
+We formalize the synchronization invariants (position IDs, attention masks, KV-cache alignment) that any valid implementation must satisfy and show precisely how prior methods fail to maintain them, producing corrupted outputs.
+  • We present EqSpec, the first batch speculative decoding algorithm that guarantees algorithmic correctness through explicit synchronization, with formal proofs that the synchronization invariants are preserved.
+  • We introduce EXSpec, which reduces alignment overhead through cross-batch scheduling that dynamically groups same-length sequences.
+On SpecBench across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B pairs, our methods achieve up to $3\times$ throughput at batch size 8 while maintaining algorithmic correctness.
 2 Design Space Analysis
@@
 We identify three approaches to handle raggedness: Masking, Rollback, and Dynamic Padding.
-Yet, as our systematic analysis shows, current instantiations of these approaches fail to simultaneously maintain correctness and performance at scale.
-To close this gap, we first analyze the pitfalls of each approach and then introduce a correctness-first algorithmic design with explicit synchronization requirements for reliable batch speculative decoding.
+Yet, as our systematic analysis shows, current instantiations of these approaches fail to simultaneously maintain algorithmic correctness and performance at scale.
+To close this gap, we first analyze the pitfalls of each approach and then introduce the first valid batch speculative decoding with explicit synchronization requirements.
 ✗Masking Approach (non-contiguous position IDs).
@@
 https://github.com/SafeAILab/EAGLE/issues/250 (Li et al., 2025) encounters similar framework limitations.
-Supporting non-contiguous position IDs would require custom CUDA kernels for each base model (Qian et al., 2024)—a prohibitive engineering cost that sacrifices portability.
+BASS (Qian et al., 2024) addresses this through custom attention kernels that tolerate ragged padding in the middle and right, but this requires coupled modifications to both attention computation and KV-cache indexing, sacrificing portability across model architectures and attention variants.
 ✗ Rollback Approach (speculation waste).
 After each verification step, all sequences are truncated to the batch’s minimum accepted length (Wolf et al., 2020; kamilakesbi, 2024).
-This guarantees alignment but discards correctly verified tokens from faster sequences.
+This keeps the batch aligned, but it throws away additional verified tokens that faster sequences had already accepted.
 As batch size grows and acceptance-rate variance widens, the waste compounds; in the extreme, one persistently rejecting sequence forces single-token progress for the entire batch.
 In effect, throughput collapses to that of the worst-performing sequence, undermining speculative gains and rendering the approach impractical at larger scales.
-Algorithm 1 BatchVerify: Single Forward Pass
-1: Target model $\mathcal{M}_{t}$, Sequences $\mathcal{S}$, Draft tokens $D$, KV cache
-2: Accepted tokens AA, Bonus tokens $B$
-3: if first iteration then
-4:   $\mathcal{X}\leftarrow\mathcal{S}\oplus D$
-5: else
-6:   $\mathcal{X}\leftarrow D$
-7: $\textit{logits},\textit{KVCache}\leftarrow\mathcal{M}_{t}(\mathcal{X},\textit{KVCache})$
-8: $\textit{pred\_tokens}\leftarrow\arg\max(\textit{logits, dim=vocab})$
-9: $\triangleright$ Vectorized first mismatch detection
-10: $\textit{matches}\leftarrow(\textit{pred\_tokens}=D)$
-11: $J\leftarrow\arg\max(\neg\textit{matches, dim=seq})$
-12: $\triangleright$ Ragged shape acceptance, no vectorization
-13: for each sequence $i$ in batch do
-14:   $A[i]\leftarrow D[i][:j]$
-15:   $\triangleright$ Get bonus token from first mismatch
-16:   $\textit{bonus\_logit}\leftarrow\textit{logits}[i,|\mathcal{S}[i]|+j]$
-17:   $B[i]\leftarrow\arg\max(\textit{bonus\_logit})$
-18: return $A,B,KVCache$
-Algorithm 2 EqSpec
-1: Draft model $\mathcal{M}_{d}$, Target model $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Max tokens $T$, Draft length $K$
-2: Generated sequences $\mathcal{S}$
-3: $\mathcal{S}\leftarrow\texttt{Tokenize}(\mathcal{P})$ $\triangleright$ Batch left padding
-4: $\textit{KVCache}\leftarrow\emptyset$
-5: while until max new tokens do
-6:   Phase 1: Draft Generation
-7:   $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{S},K)$
-8:   Phase 2: Batch Verification
-9:   $A,B,\textit{KVCache}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{S},D,\textit{KVCache})$
-10:   $\triangleright$ See Figure 3 for illustration on index offset.
-11:   Phase 3: Unpad-Append-Repad
-12:   for each sequence $i$ in batch do
-13:    $\mathcal{S}[i]\leftarrow\texttt{Unpad}(\mathcal{S}[i])$
-14:    $\mathcal{S}[i]\leftarrow\mathcal{S}[i]\oplus A[i]\oplus B[i]$
+A variant approach, employed by vLLM’s v0 engine, avoids truncation through batch expansion: each sequence is duplicated $K$ times with progressively longer draft prefixes (e.g., $t_{1}],[t_{1},t_{2}],\ldots$), all verified in a single forward pass.
+While this preserves accepted tokens, it incurs $K\times$ redundant computation and memory overhead that scales with both batch size and draft length.
+At scale, this memory explosion caused GPU out-of-memory failures, ultimately leading to the deprecation of batch expansion in vLLM v1 (see Appendix C for details).
+✓ Dynamic Padding Approach.
+This approach realigns sequences after each verification by adjusting left padding to maintain right alignment, preserving all accepted tokens.
+While conceptually simple, algorithmic correctness requires tight synchronization of position IDs, attention masks, and the KV-cache.
+DSD’s experimental code (Yan et al., 2025) follows this idea but merely repads at each step—adding varying left padding without ever unpadding—thereby inflating sequences.
+It also contains three critical errors: (i) sampling bonus tokens from the draft model rather than the target model; (ii) redundantly regenerating KV-cache entries, causing memory bloat; and (iii) desynchronizing padding, position IDs, and the KV-cache across iterations.
+Despite the overhead of repeated realignment, a valid dynamic-padding implementation fits within standard frameworks and preserves all verified tokens.
+Among the three approaches, only dynamic padding is viable.
+(1) Masking with non-contiguous position IDs addresses only half the problem: even if attention is patched to tolerate such layouts (via adjusted position IDs and masking), algorithmic correctness still requires that the KV-cache correspond to the same post-verification token stream across verification rounds.
+If one fully fixes both sides, the implementation must perform per-round cache/token realignment (i.e., compacting/realigning the batch state after verification), which effectively converges to dynamic padding.
+(2) Rollback preserves alignment but discards verified tokens by truncating all sequences to the batch’s minimum accepted length, and this waste grows with batch size.
+(3) Dynamic padding maintains algorithmic correctness by explicitly synchronizing both position IDs and KV-cache after each verification round.
+We provide a detailed bug taxonomy in Appendix A, distinguishing batch-independent implementation errors (e.g., wrong sampling distribution) from ragged tensor synchronization failures that only manifest at batch size >1>1.
+We identify and address all of these, but the synchronization problem is the core challenge: no prior work formalizes what invariants batch speculative decoding must maintain, and this formalization is our central contribution.
+Figure 3: EqSpec synchronizes via unpad–append–repad.
+Output Equivalence.
+We formalize the validity requirement for batch speculative decoding.
+Under greedy decoding (temperature=0{}=0), speculative decoding must yield outputs identical to standard autoregressive generation, and any token-level divergence indicates an implementation error.
+For temperature>0{}>0, where outputs are sampled stochastically, speculative decoding must preserve the output distribution of the target model; that is, the probability of generating any token sequence must remain unchanged.
+This losslessness, whether measured by exact output equivalence or distributional equivalence, is the defining criterion of valid speculative decoding and the hallmark of an authentic implementation.
+Methods that achieve high throughput while violating this criterion, as we demonstrate for BSP and DSD in Figure.1, do not constitute valid speculative decoding regardless of their reported performance metrics.
+3 Method
+We propose a batch speculative decoding method that explicitly maintains synchronization across sequences (position IDs, attention masks, and KV-cache) while also aiming to keep the added synchronization overhead small for higher throughput.
+The main trade-off is that maintaining equivalence to standard decoding requires re-synchronizing position IDs, attention masks, and the KV-cache after each verification round when the batch becomes ragged, and this synchronization can account for up to 40% of total computation.
+We introduce two complementary mechanisms: EqSpec specifies and enforces the minimal synchronization invariants required for valid computation (Section 3.1), while EXSpec groups same-length sequences to avoid synchronization overhead (Section 3.2).
+3.1 EqSpec: Algorithmic Correctness via Synchronization Invariants
+Rectangular Alignment.
+We formalize the algorithmic correctness requirements for batch speculative decoding.
+Let $B$ denote the batch size, and let sequence $i$ have padding length $p_{i}$, content length $c_{i}$, and $a_{i}$ newly accepted tokens (including the bonus token) after verification round $t$.
+A valid implementation must maintain rectangular alignment across the batch:
+$\quad\forall i\in[1,B],\quad p_{i}^{(t)}+c_{i}^{(t)}=L^{(t)}$ (I1)
+where $L^{(t)}$ is the padded sequence length at round $t$.
+Position-ID Contiguity.
+Additionally, position IDs must be derived consistently from the attention mask:
+$\quad\mathrm{pos}_{i}[j]=\max\left(0,\sum_{k=0}^{j}\mathbf{1}[\mathrm{tok}_{i}[k]\neq\texttt{pad}]-1\right)$ (I2)
+which guarantees contiguous position IDs starting from zero at the first content token, as required by RoPE (Su et al., 2024) and other position-aware attention mechanisms.
+Continuous batching systems (e.g., vLLM, SGLang) process requests using variable-length packing, sidestepping I1, but still require scatter-gather across requests and rollback both position-ID and KV-cache for rejected tokens—a one-dimensional analog of alignment.
+Critically, I2 remains a prerequisite for algorithmic correctness regardless of batching strategy; note that neither vLLM nor SGLang currently supports continuous batching speculative decoding with external draft models (see Appendix.E).
+After batch verification, sequence $i$ accepts $a_{i}$ tokens, breaking rectangular alignment since $a_{i}$ varies across sequences. To restore I1, the new padding must satisfy:
+$\forall i\in[1,B],\quad p_{i}^{(t+1)}+c_{i}^{(t)}+a_{i}=L^{(t+1)}$ (1)
+where $L^{(t+1)}=\max_{j}(c_{j}^{(t)}+a_{j})$. This yields the per-sequence padding offset:
+$\delta_{i}=p_{i}^{(t+1)}-p_{i}^{(t)}=\left(L^{(t+1)}-L^{(t)}\right)-a_{i}$ (2)
+We formally prove that this update rule preserves rectangular alignment I1 across all verification rounds (Theorem B.1 in Appendix).
+Invariant I2 is enforced by recomputing position IDs after each repadding operation.
+Algorithm 1 EqSpec: Minimalism Batch Spec
+0:  Draft model $\mathcal{M}_{d}$, Target model $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Max tokens $T$, Draft length $K$
+0:  Generated sequences $\mathcal{S}$
+1:  $\mathcal{S}\leftarrow\texttt{Tokenize}(\mathcal{P})$ $\triangleright$ Batch left padding
+2:  $\textit{KVCache}\leftarrow\emptyset$
+3:  while until max new tokens do
+4:  Phase 1: Draft Generation
+5:  $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{S},K)$
+6:  Phase 2: Batch Verification
+7:  $\triangleright$ See Alg. 3 in Appendix for BatchVerify
+8:  $A,B,\textit{KVCache}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{S},D,\textit{KVCache})$
+9:  $\triangleright$ See Figure 3 for illustration on index offset.
+10:  Phase 3: Unpad-Append-Repad
+11:  for each sequence $i$ in batch do
+12:   $\mathcal{S}[i]\leftarrow\texttt{Unpad}(\mathcal{S}[i])$
+13:   $\mathcal{S}[i]\leftarrow\mathcal{S}[i]\oplus A[i]\oplus B[i]$
+14:  end for
 15:  $\mathcal{S},\textit{offset}\leftarrow\texttt{BatchRepad}(\mathcal{S})$
 16:  $\textit{KVCache}\leftarrow\texttt{Realign}(\textit{KVCache},\textit{offset})$
-17: return $\mathcal{S}$
-✓ Dynamic Padding Approach.
-This approach realigns sequences after each verification by adjusting left padding to maintain right alignment, preserving all accepted tokens.
-While conceptually simple, correctness requires tight synchronization of position IDs, attention masks, and the KV-cache.
-DSD’s experimental code (Yan et al., 2025) follows this idea but merely repads at each step—adding varying left padding without ever unpadding—thereby inflating sequences.
-It also contains three critical errors: (i) sampling bonus tokens from the draft model rather than the target model; (ii) redundantly regenerating KV-cache entries, causing memory bloat; and (iii) desynchronizing padding, position IDs, and the KV-cache across iterations.
-Despite the overhead of repeated realignment, a correct dynamic-padding implementation fits within standard frameworks and preserves all verified tokens.
-Among the three approaches, only dynamic padding is viable: position-ID schemes require custom kernels that undermine portability, and rollback discards verified tokens at rates that grow with batch size; dynamic padding maintains correctness within standard frameworks.
-3 Method
-We present a synchronization-aware approach to batch speculative decoding that co-designs correctness and efficiency.
-The core tension is that preserving correctness requires synchronizing position IDs, attention masks, and the KV-cache across ragged tensors—an overhead that can consume 40% of computation.
-We introduce two complementary mechanisms: EqSpec specifies and enforces the minimal synchronization invariants required for valid computation (Section 3.1), while EXSpec groups same-length sequences to avoid synchronization overhead (Section 3.2).
-Together, these form a unified system in which correctness constraints drive scheduling, enabling both output equivalence and practical performance.
-Figure 3: EqSpec synchronizes via unpad–append–repad.
-3.1 Minimal Batch Realignment: EqSpec
-Figure 3 illustrates the core challenge and our remedy for maintaining correctness in batch speculative decoding.
-After each verification round, sequences accept different numbers of draft tokens, producing ragged tensors that GPUs cannot process.
-To restore a valid batch, we apply an unpad–append–repad procedure that converts ragged outputs back to a rectangular layout while preserving three invariants: contiguous position IDs, valid attention masks, and aligned KV-cache entries.
-Correct implementation requires padding-agnostic position IDs that reset to zero at the first content token, attention masks that exclude padding tokens, and precise KV-cache realignment after each verification step (Algorithm 1).
-After unpad–append–repad, index offsets shift across sequences; consequently, position IDs and attention masks must be recomputed to preserve correct token relationships.
-Moreover, the bonus token introduces a special case: it is sampled from the target model’s output distribution at the first mismatch position after the verification forward pass has already completed.
-Consequently, while it encodes the target’s authoritative correction, it lacks KV-cache entries in either the draft or target model because KV-cache is computed only for tokens that were present in the input sequence during the forward pass—and the bonus token, being an output, was not yet part of that input sequence.
-Therefore, it must be included in the next forward pass to create its KV-cache entries, further complicating synchronization.
-Finally, the realignment is resource-intensive: the KV-cache consists of rank-4 tensors $(\text{batch}\times\text{heads}\times\text{sequence}\times\text{dimension})$, and each padding adjustment triggers allocation and concatenation of high-dimensional zeros.
-Algorithm 2 details the complete EqSpec procedure.
+17:  end while
+18:  return $\mathcal{S}$
+Unpad-Append-Repad.
+Figure 3 illustrates how EqSpec restores a valid batch after each verification round.
+The procedure applies the offset $\delta_{i}$ to realign all batch state: token sequences receive fresh left padding, attention masks are recomputed, position IDs are recalculated via I2, and KV-cache entries are shifted by $\delta_{i}$ positions to maintain token-to-cache correspondence.
+During batch verification, the target model performs a single forward pass over draft tokens, detects the first mismatch per sequence, and samples a bonus token from the target distribution (Algorithm 3 in Appendix.C).
+However, since this sampling occurs after the forward pass completes, the bonus token lacks KV-cache entries and must be included in the next forward pass to create them.
+The complete EqSpec procedure is detailed in Algorithm 1.
 Figure 4: EXSpec pools ragged sequences by length, avoiding realignment; only unmatched sequences need syncing, turning fixed overhead into optional cost.
-Speedup Analysis.
+Algorithm 2 EXSpec: Cross-Batch Scheduling
+0:  Draft and Target model $\mathcal{M}_{d}$, $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Window size $W$, Batch size $B$
+0:  Generated sequences $\mathcal{S}$
+1:  $\textit{Pool}\leftarrow\texttt{InitSequencePool}(\mathcal{P})$
+2:  $\triangleright$ Tokenize and optionally sort by length
+3:  $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
+4:  while Pool.hasActive() do
+5:  Phase 1: Lazy Realignment
+6:  $\mathcal{B},\textit{mask},\textit{KV}\leftarrow\texttt{GetBatch}(\textit{Window},B)$
+7:  $\triangleright$ Try same-length concatenation
+8:  $\triangleright$ Fallback to Unpad-Repad Realignment
+9:  Phase 2: Draft Generation
+10:  $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{B},\textit{mask},K)$
+11:  Phase 3: Batch Verification
+12:  $\triangleright$ See Alg. 3 in Appendix for BatchVerify
+13:  $A,B,\textit{KV}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{B},D,\textit{KV})$
+14:  Phase 4: Write-Back and Window Refill
+15:  for $i\in\mathcal{B}$ do
+16:   $\textit{Pool}[i]\leftarrow\textit{Pool}[i]\oplus A[i]\oplus B[i]$
+17:   $\textit{Pool.KV}[i]\leftarrow\textit{KV}[i]$
+18:   if $\textit{isComplete}(\textit{Pool}[i])$ then $\textit{Pool.deactivate}(i)$
+19:  end for
+20:  $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
+21:  end while
+22:  return Pool.sequences
+3.2 EXSpec: Reducing Overhead via Cross-Batch Scheduling
+Theoretical Speedup Analysis
 The speedup of speculative decoding depends on both the token-acceptance rate and computational costs.
@@
 (2023) analyzes single-sequence performance, modeling the expected tokens generated per iteration as $(1-\alpha^{k+1})/(1-\alpha)$, where $\alpha$ is the token acceptance rate (TAR) and $k$ is the number of draft tokens per speculation round.
-This single-sequence view assumes negligible batch overhead and focuses purely on acceptance dynamics.
-Batch speculative decoding, however, introduces additional complexities not captured by that model—most notably alignment overhead, which can dominate and degrade performance.
-We therefore introduce a batch-aware speedup model:
+This single-sequence view assumes no batch overhead and focuses purely on acceptance dynamics.
+For batch speculative decoding, we introduce a batch-aware speedup formula:
 $S\;=\;\frac{\alpha\cdot k}{\,c_{\text{draft}}+c_{\text{verify}}+c_{\text{overhead}}(B)\,}$ (3)
 where $S$ denotes speedup relative to non-speculative decoding (i.e., running the target model alone), $c_{\text{draft}}$ and $c_{\text{verify}}$ are the relative costs of draft generation and target verification, and $c_{\text{overhead}}(B)$ captures batch-dependent alignment overhead absent from single-sequence analysis.
-Algorithm 3 EXSpec: Cross-Batch Scheduling
-1: Draft and Target model $\mathcal{M}_{d}$, $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Window size $W$, Batch size $B$
-2: Generated sequences $\mathcal{S}$
-3: $\textit{Pool}\leftarrow\texttt{InitSequencePool}(\mathcal{P})$
-4: $\triangleright$ Tokenize and optionally sort by length
-5: $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
-6: while Pool.hasActive() do
-7:   Phase 1: Lazy Realignment
-8:   $\mathcal{B},\textit{mask},\textit{KV}\leftarrow\texttt{GetBatch}(\textit{Window},B)$
-9:   $\triangleright$ Try same-length concatenation
-10:   $\triangleright$ Fallback to Unpad-Repad Realignment
-11:   Phase 2: Draft Generation
-12:   $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{B},\textit{mask},K)$
-13:   Phase 3: Batch Verification
-14:   $A,B,\textit{KV}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{B},D,\textit{KV})$
-15:   Phase 4: Write-Back and Window Refill
-16:   for $i\in\mathcal{B}$ do
-17:    $\textit{Pool}[i]\leftarrow\textit{Pool}[i]\oplus A[i]\oplus B[i]$
-18:    $\textit{Pool.KV}[i]\leftarrow\textit{KV}[i]$
-19:    if $\textit{isComplete}(\textit{Pool}[i])$ then $\textit{Pool.deactivate}(i)$
-20:   $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
-21: return Pool.sequences
-Two observations follow.
-First, when sequences within a batch share the same length, batch realignment overhead is negligible; likewise, with very small draft models or prompt lookahead decoding (Fu et al., 2024b), $c_{\text{draft}}$ becomes small relative to verification.
-Second, and critically, $c_{\text{overhead}}(B)$ scales superlinearly with batch size $B$ due to the ragged-tensor problem (Section 4.3).
-This overhead decomposes as $c_{\text{overhead}}(B)=c_{\text{pad}}(B)+c_{\text{kv}}(B)$, where $c_{\text{pad}}(B)$ accounts for unpad–append–repad operations and $c_{\text{kv}}(B)$ for KV-cache realignment on rank-4 tensors.
-While $c_{\text{draft}}$ and $c_{\text{verify}}$ benefit from GPU parallelism and remain relatively stable within the hardware’s batch-parallel regime, the alignment overhead grows with both batch size and variance in acceptance rates across sequences.
-3.2 Cross-Batch Scheduling: EXSpec
-Profiling EqSpec shows that alignment overhead consumes 39.4% of computation at batch size 8, rising to 46.7% at batch size 16.
-Because this cost is inherent to synchronizing ragged tensors, micro-optimizing the primitives yields limited gains.
-Instead of accelerating these operations, EXSpec avoids them by scheduling.
+Critically, $c_{\text{overhead}}(B)$ scales superlinearly with batch size $B$.
+Inside the alignment overhead, the KV-cache is the major bottleneck, consisting of rank-4 tensors (batch $\times$ heads $\times$ sequence $\times$ dimension), and each padding adjustment triggers allocation and concatenation of high-dimensional zeros.
+While $c_{\text{draft}}$ and $c_{\text{verify}}$ benefit from GPU parallelism, the alignment overhead grows with both batch size and variance in acceptance rates across sequences.
+Profiling EXSpec confirms this: alignment consumes 39.4% of computation at batch size 8, rising to 46.7% at batch size 16.
+However, the speedup analysis reveals a key insight: when all sequences accept the same number of tokens, $\delta_{i}=0$ and realignment overhead vanishes (Corollary B.3).
+Rather than accelerating these operations, EXSpec sidesteps them through scheduling.
+Cross-Batch Scheduling.
 EXSpec differs from EqSpec in how it manages sequence lifecycles.
 Rather than maintaining fixed batches that require realignment after every verification, we introduce a SequencePool that holds sequences individually in their ragged states.
-This enables three optimizations: (i) sequences that complete (reach EOS) are immediately removed, avoiding wasted computation on finished items; (ii) lazy realignment defers synchronization until strictly necessary, keeping sequences in their natural ragged form between steps; and (iii) dynamic batch formation over a sliding window of $W>B$ sequences greatly increases the chance of finding same-length groups that can be concatenated without any realignment.
+This enables three optimizations: (i) sequences that complete (reach EOS) are immediately removed, avoiding wasted computation; (ii) lazy realignment defers synchronization until strictly necessary; and (iii) dynamic batch formation over a sliding window of $W>B$ sequences greatly increases the chance of finding same-length groups.
 Figure 4 illustrates the flow.
 After verification, ragged sequences return directly to the pool without realignment.
-The scheduler then scans the active window and attempts to form batches of identical length.
-Same-length sequences concatenate directly—no padding adjustments, no position-ID recomputation, no KV-cache realignment—bypassing $c_{\text{overhead}}(B)$.
-Only when same-length grouping fails do we fall back to the expensive unpad–append–repad procedure from EqSpec.
-Algorithm 3 formalizes this in four phases: dynamic batch formation (prioritizing same-length groups), draft generation, verification, and pool write-back with continuous window refresh.
-Combined with prompt-length sorting, grouping rates approach unity for similar workloads, turning the worst case (constant realignment) into the rare case.
+The scheduler scans the active window and attempts to form batches of identical length.
+Same-length sequences concatenate directly—no padding adjustments, no position-ID recomputation, no KV-cache realignment—bypassing $c_{\text{overhead}}(B)$ entirely.
+Only when same-length grouping fails do we fall back to unpad–append–repad.
+Combined with prompt-length sorting, grouping rates approach unity for similar workloads, turning constant realignment into the rare case.
+The complete procedure is detailed in Algorithm 2.
 4 Experiments
-We evaluate EqSpec and EXSpec along two dimensions often conflated in prior work: correctness (whether outputs match non-speculative generation) and throughput (tokens per second).
+We evaluate EqSpec and EXSpec along two dimensions often conflated in prior work: output equivalence and throughput.
 Through systematic evaluation across three model families and comparisons with both research prototypes and production systems, we show that our approach uniquely preserves output equivalence while achieving competitive speedups.
@@
 Evaluation and Datasets.
-We use SpecBench (Xia et al., 2024), focusing on the first turn of each conversation because our evaluation targets offline batch inference rather than multi-turn interaction.
+We use SpecBench (Xia et al., 2024), a widely adopted benchmark for speculative decoding research.
 We also use Multi30k (Elliott et al., 2016) for a controlled EXSpec study that contrasts random sampling with an identical-length subset, isolating sequence-length diversity as the driver of grouping rate.
-For the main evaluation, we measure: (1) Throughput: tokens/s across batch sizes; and (2) Correctness: exact-match rate (full-sequence equivalence with non-speculative decoding) and partial-match rate (fraction of tokens matching until the first divergence).
+For the main evaluation, we measure: (1) Throughput: tokens/s across batch sizes; and (2) Output Equivalence: exact-match rate (full-sequence equivalence with non-speculative decoding) and partial-match rate (fraction of tokens matching until the first divergence).
 The partial-match metric helps localize failure modes—early divergence typically indicates position-ID or KV-cache misalignment.
@@
 Following our design-space taxonomy (Section 2), we evaluate: (1) Masking approaches: BSP (Su et al., 2023) attempts masking with adaptive speculation but suffers position-ID inconsistencies (BASS (Qian et al., 2024) also follows this approach but requires custom CUDA kernels, limiting generality); (2) Dynamic-padding approaches: DSD (Yan et al., 2025) explores padding but mishandles the KV-cache, while EqSpec implements correct synchronization and EXSpec adds cross-batch scheduling; (3) Reference baselines: Spec-1 (batch-size-1 speculation from Hugging Face Transformers), which does not support batch speculative decoding [^2] https://github.com/huggingface/transformers/issues/32165.
-We also compare with production systems vLLM [^3] We use the vLLM v0 engine because v1 deprecates speculative decoding.
-(Kwon et al., 2023) (which subsumes TETRIS (Wu et al., 2025) and TurboSpec (Liu et al., 2025) as vLLM forks) and SGLang-EAGLE [^4] SGLang is compatible only with the EAGLE family; we compare Vicuna-7B/EAGLE2 and Qwen3-8B/EAGLE3.
-There are no available weights for GLM-4.
-(Zheng et al., 2024; Li et al., 2024b; 2025), noting that their continuous batching and memory-management designs complicate direct comparison.
-We further test SGLang-EAGLE-Deterministic (SGLang Team, 2025), which enables deterministic execution to reduce numerical variance.
+Production systems (vLLM, SGLang; see also Sections 2 and 3.1) employ continuous batching with paged attention, which sidesteps the ragged tensor problem through different architectural choices.
+However, neither currently supports batch speculative decoding with external draft models: vLLM deprecated this in v1 due to batch expansion memory overhead, and SGLang only supports EAGLE-family drafters.
+We include a comparison in Appendix E for completeness, but note these systems solve a related but distinct problem.
 No existing implementation uses rollback due to its inherent wastefulness.
@@
 E P E P E P E P E P E P
-Batch-based Methods (vs. batch=1 non-spec)
 Non-Spec-Batch 53.8 98.2 92.9 96.5 93.3 97.2
@@
 BSP 1.9 39.7 0.2 31.3 3.5 19.9 2.1 12.6 1.0 15.3 0.6 8.1
-Continuous Batching Systems (vs. own non-spec)
-vLLM + Spec 96.9 / 98.0 65.6 / 78.5 72.7 / 84.5
-SGLang + EAGLE 69.8 / 79.5 47.7 / 65.4
-SGLang + EAGLE + Det 85.0 / 90.0 50.6 / 69.5
-Table 1: Correctness check reports exact match (E) and partial match (P) scores.
-Top: compares with batch=1 non-spec baseline.
-Bottom: with each system’s non-speculative mode (scores reported as E / P).
+Table 1: Output Equivalence of speculative decoding implementations.
+Exact match (E) and partial match (P) scores compared against batch=1 non-speculative baseline.
 Our approach sustains $95\%$ accuracy, while prior work (DSD, BSP) suffers major drops from KV-cache and position ID errors.
 Figure 5: Decomposing batch speculative decoding performance.
-(a) Batch scaling efficiency normalized to BS=1, isolating GPU parallelism from per-sequence speculation; EXSpec initially exceeds the No-Ragged-Scaling upper bound before degrading due to alignment overhead.
+(a) Batch scaling efficiency: each method’s throughput normalized to its own BS=1 baseline, isolating scaling behavior from absolute performance.
+The No-Ragged-Scaling line shows measured autoregressive decoding throughput (not a theoretical bound), serving as a reference for how standard batching scales without ragged tensor overhead.
 (b) Alignment overhead grows super-linearly with batch size, consuming up to 38% of inference time, validating that $c_{\text{overhead}}(B)$ dominates at scale.
@@
 uniform-length sequences, showing that length homogeneity transforms grouping effectiveness.
-4.2 Output Correctness Verification
-We verify correctness using deterministic greedy decoding to eliminate sampling variance and enable precise bug isolation.
+4.2 Output Equivalence Verification
+We verify output equivalence using deterministic greedy decoding to eliminate sampling variance and enable precise bug isolation.
 This avoids metrics such as ROUGE (Qian et al., 2024), which can mask implementation failures (e.g., repetitive corruption can still score reasonably).
@@
 Table 1 reveals distinct patterns.
-Our methods maintain $\approx$95% exact match across settings, the remaining $\approx$5% divergence stems from numerical non-determinism in floating-point operations (He & Lab, 2025) and tie-breaking in argmax sampling rather than algorithmic errors.
+Our methods maintain $\approx$95% exact match across settings, the remaining $\approx$5% divergence stems from numerical non-determinism in floating-point operations (He and Lab, 2025; Gond et al., 2026) and tie-breaking in argmax sampling rather than algorithmic errors.
+This attribution is corroborated by Liu et al.
+(2025b), who observe that generation length varies significantly across batch sizes due to kernel-level nondeterminism even in functionally correct implementations.
 By contrast DSD and BSP fail catastrophically with different signatures.
@@
 These complementary patterns—immediate failure vs.
-gradual decay—show how partial-match metrics reveal not just that implementations fail, but when and why.
-Production systems (vLLM, SGLang) introduce additional complexity.
-While both are accurate on Vicuna, they degrade markedly on Qwen3.
-SGLang-EAGLE-Deterministic helps disambiguate the cause: improved accuracy suggests most divergences stem from floating-point non-determinism (He & Lab, 2025) rather than algorithmic bugs.
-Despite this, we show below that their throughput still falls below non-speculative baselines.
+gradual decay—reflect distinct root causes: DSD suffers from batch-independent errors (wrong sampling distribution) compounded by synchronization failures, while BSP accumulates KV-cache drift from position-ID desynchronization.
+A complete taxonomy distinguishing batch-independent bugs from batch-specific synchronization errors appears in Appendix A.
+We provide additional analysis of production systems (vLLM, SGLang) in Appendix E.
 4.3 Overhead and Scaling Dynamics
@@
 Figure 1 shows that EXSpec outperforms EqSpec in absolute throughput by combining speculative and batching gains, yet EqSpec exhibits negative scaling beyond BS=8.
-To test whether batch speculative decoding can retain GPU-parallelism benefits despite raggedness, Figure 5(a) normalizes each method’s throughput to its BS=1 baseline, isolating batch-level scaling from per-sequence speculation and enabling comparison to No-Ragged-Scaling (standard batched decoding without speculation).
-A notable effect emerges: EXSpec initially surpasses even this upper bound because speculation converts memory-bound operations into compute-bound verification, improving GPU utilization when managed correctly.
+To test whether batch speculative decoding can retain GPU-parallelism benefits despite raggedness, Figure 5(a) measures batch scaling efficiency: each method’s throughput at batch size $N$ divided by its own throughput at batch size 1.
+This normalization isolates how well each method scales with batch size, independent of absolute performance differences.
+The No-Ragged-Scaling line represents actual measured autoregressive decoding throughput (without speculation), similarly normalized to its own BS=1 baseline; it is not a theoretical upper bound but rather a reference showing how standard batching scales when no ragged tensor synchronization is required.
+A notable effect emerges: EXSpec initially exceeds this reference line.
+This occurs because each method is normalized independently, and speculation converts memory-bound token generation into compute-bound verification, which benefits more from GPU parallelism.
+Although EXSpec may have lower absolute throughput than standard decoding at BS=1 due to draft model overhead, its relative scaling from BS=1 to BS=8 can surpass that of non-speculative decoding.
 However, at larger batch sizes, alignment overhead dominates and the advantage inverts, confirming that $c_{\text{overhead}}(B)$ eventually overwhelms parallelism gains.
-Figure 6: Speculative decoding lags non-speculative baselines, with larger batches further degrading throughput.
+The key finding is that valid batch speculative decoding need not sacrifice batch scaling efficiency, given that prior methods either produce gibberish outputs or fail to scale.
 Alignment overhead growth.
@@
 This contrast suggests that preprocessing strategies (e.g., bucketing, dynamic sorting) can push real workloads toward the ideal, revealing untapped potential when scheduling is paired with workload shaping.
-Production systems limitation.
-Figure 6 shows that production frameworks struggle with batch speculative decoding.
-Although our method is not directly comparable to continuous-batching systems (which vary effective batch size dynamically), the trend is consistent: vLLM’s speculative decoding underperforms its baseline at high concurrency (leading to deprecation in v1), and SGLang+EAGLE is slower than non-speculative generation across all batch sizes.
-Two architectural factors likely contribute: (i) speculation repurposes the prefill stage for parallel verification, breaking the prefill–decode separation these systems optimize; and (ii) continuous batching already induces raggedness from varying request lengths, and speculation compounds this with per-sequence acceptance variance, creating nested raggedness that overwhelms existing schedulers.
-These outcomes indicate that speculation cannot simply be grafted onto architectures optimized for predictable token-by-token decoding; deeper redesign is required.
-Method TPS $|\text{Spec}|$ Time/Verif. (ms) Verif. % Draft % Overhead %
-EqSpec 95.6 1469 24.88 44.9 27.4 27.7
-EXSpec 156.4 952 29.13 55.9 29.5 14.6
+Method TPS $|\text{Spec}|$ ms/V V% D% OH%
+EqSpec 95.6 1469 24.9 44.9 27.4 27.7
+EXSpec 156.4 952 29.1 55.9 29.5 14.6
 Table 2: Overhead anatomy of batch speculation methods.
-Despite slower per-verification time, EXSpec achieves higher throughput by reducing total verification calls and minimizing alignment overhead through intelligent scheduling.
+V%=verification, D%=draft, OH%=overhead.
+Despite slower per-verification time, EXSpec achieves higher throughput by reducing verification calls and alignment overhead.
 Overhead Anatomy.
@@
 Speculative decoding accelerates LLM inference by verifying draft tokens in parallel.
-Two verification paradigms dominate: sequence verification—as in SpecDec and follow-ups (Xia et al., 2023; Santilli et al., 2023; Yang et al., 2023; Hooper et al., 2023; Zhang et al., 2023; Fu et al., 2024a)—which preserves the target model’s output distribution; and tree verification—e.g., Medusa/EAGLE variants (Miao et al., 2024; Spector & Re, 2023; Sun et al., 2023; He et al., 2023; Cai et al., 2024; Li et al., 2024a; b; 2025)—which explores multiple branches to raise acceptance rates.
+Two verification paradigms dominate: sequence verification—as in SpecDec and follow-ups (Xia et al., 2023; Santilli et al., 2023; Yang et al., 2023; Hooper et al., 2023; Zhang et al., 2023; Fu et al., 2024)—which preserves the target model’s output distribution; and tree verification—e.g., Medusa/EAGLE variants (Miao et al., 2024; Spector and Re, 2023; Sun et al., 2023; He et al., 2023; Cai et al., 2024; Li et al., 2024b, a, 2025)—which explores multiple branches to raise acceptance rates.
 Recent works parallelize multiple draft sequences per request (Stewart et al., 2024; Lee et al., 2024) but still verify each request independently.
@@
 Position ID/masking approaches like BSP (Su et al., 2023) suffer from position ID inconsistencies across iterations.
-Dynamic padding approaches reveal critical implementation errors: both DSD (Yan et al., 2025) and Meta’s recent work (Tang et al., 2025) incorrectly sample bonus tokens from the draft model’s distribution rather than the target model’s, violating the fundamental correctness guarantee of speculative decoding.
+Dynamic padding approaches reveal critical implementation errors: both DSD (Yan et al., 2025) and Meta’s recent work (Tang et al., 2025) incorrectly sample bonus tokens from the draft model’s distribution rather than the target model’s, violating the fundamental output equivalence guarantee of speculative decoding.
 This error compounds with improper KV-cache handling, producing corrupted outputs—BSP generates repetitive tokens while DSD produces <unk> symbols.
 BASS (Qian et al., 2024) sidesteps these issues through custom CUDA kernels but sacrifices portability.
-These failures demonstrate that the ragged tensor problem extends beyond performance to correctness itself, motivating our cross-batch scheduling approach that maintains correctness while achieving batch scaling without custom kernels.
-Production Systems for Speculative Decoding.
-Serving stacks integrate speculation atop general batching (vLLM, SGLang/EAGLE, and forks such as TETRIS and TurboSpec (Kwon et al., 2023; Zheng et al., 2024; Li et al., 2025; Wu et al., 2025; Liu et al., 2025)).
-In practice, variable draft acceptance clashes with continuous batching and paged attention, and our measurements indicate scaling limits (e.g., small single-sequence gains that reverse at higher concurrency); some engines have since reduced or deprecated speculative-decoding support.
-We intentionally build our reference implementation from scratch, prioritizing correctness checks; compatibility with continuous batching, paged attention, and prefill–decode separation is orthogonal and left as future work.
+Production systems (vLLM, SGLang) avoid the problem entirely via continuous batching but do not support external draft models at batch scale (Appendix E).
+These failures reveal that batch speculative decoding is fundamentally challenge—yet no prior work formalizes what invariants must hold.
+We address this gap by establishing synchronization requirements and providing the first correct implementation.
 6 Conclusion
-We presented a correctness-first approach to batch speculative decoding that resolves the fundamental tension between maintaining output equivalence and achieving practical speedups.
-By identifying the precise synchronization invariants required for correctness, we explained why existing methods fail and established the minimal requirements for valid computation.
-In particular, our EqSpec implementation enforces these invariants but reveals that realignment overhead consumes up to 40% of computation.
-By contrast, EXSpec overcomes this limitation via cross-batch scheduling that dynamically groups same-length sequences to bypass realignment entirely.
-Moreover, experiments across three model families show that our approach uniquely preserves >>95% output equivalence while achieving up to 3× throughput improvement at batch size 8, hereby validating that a careful co-design of correctness constraints and scheduling can multiply batch parallelism with per-sequence speculation gains.
-Our findings on ragged shape tensor synchronization overhead and cross-batch scheduling may inform future designs for production systems, where integrating speculation with continuous batching remains an open challenge.
-Acknowledgment
-This work is supported by an eBay research grant.
-Appendix A Appendix
+We establish that valid batch speculative decoding requires maintaining synchronization invariants, specifically rectangular alignment and position-ID contiguity, which prior implementations fail to preserve.
+Our theoretical analysis formalizes these invariants and proves that the EQSPEC update rule preserves them across verification rounds, while our cost analysis reveals that alignment overhead grows superlinearly with batch size as an inherent cost of algorithmic correctness.
+EXSPEC reduces this overhead by exploiting uniform acceptance rates, where realignment cost vanishes, through cross-batch scheduling that dynamically groups same-length sequences.
+Empirically, our methods achieve 95% exact match across three model families while delivering up to 3× throughput at batch size 8, demonstrating that the synergy between batching and speculation is achievable only through principled synchronization.
+Impact Statement
+This work focuses on improving the efficiency of LLM inference without altering model outputs or introducing biases.
+By providing the first valid batch speculative decoding implementation with output equivalence guarantees, we establish a reliable foundation for future research and production deployment.
+Prior implementations produced corrupted outputs, limiting their practical utility and potentially misleading subsequent work built upon them.
+Our approach ensures that acceleration techniques do not compromise model reliability.
+We will release our implementation as open source upon acceptance, providing the research community with a verified reference implementation for batch speculative decoding.
+Appendix Contents
+  • A
+    Bug Taxonomy in Prior Implementations
+  • B
+    Proof of Synchronization Invariants
+  • C
+    BatchVerify Algorithm Details
+  • D
+    Latency-Throughput Tradeoff in Online Serving
+  • E
+    Continuous Batching Systems
+  • F
+    Discussion
+  • G
     Reproducibility Statement
+  • H
+    Disclose on LLM Usage
+Appendix A Bug Taxonomy in Prior Implementations
+We systematically categorize the implementation errors found in prior batch speculative decoding systems to clarify the distinct failure modes and their root causes.
+Table 3 distinguishes two fundamentally different bug categories based on whether they manifest at batch size 1.
+Bug Category Specific Issue System BS=1 BS>>1
+Batch-independent implementation errors Bonus token sampled from draft model instead of target model DSD
+KV cache contains rejected tokens; position IDs auto-calculated incorrectly BSP
+Ragged tensor synchronization (batch-specific) Unaligned KV cache across sequences with different accepted lengths DSD, BSP
+Position ID desynchronization across ragged sequences BSP
+Table 3: Taxonomy of implementation errors in batch speculative decoding.
+Batch-independent errors corrupt outputs at any batch size and could be fixed with careful code review.
+Ragged tensor synchronization errors are batch-specific and represent the core challenge this paper addresses.
+Batch-independent errors.
+The first category consists of relatively straightforward implementation mistakes—sampling from the wrong distribution or failing to exclude rejected tokens.
+While these corrupt outputs at any batch size, they could be fixed with careful code review.
+DSD incorrectly samples the bonus token from the draft model’s distribution rather than the target model’s, violating the fundamental output equivalence guarantee of speculative decoding.
+BSP’s KV cache retains rejected tokens and relies on framework auto-calculation for position IDs, which produces incorrect values when padding patterns change.
+Batch-specific synchronization errors.
+The second category represents the core contribution of this paper: errors that only manifest when batch size exceeds one.
+These arise from the ragged tensor problem—when sequences accept different numbers of draft tokens, maintaining synchronized position IDs, attention masks, and KV-cache state across the batch becomes non-trivial.
+Our synchronization invariants (I1, I2) in Section 3 formalize exactly what must be preserved, and our EqSpec algorithm provides the first correct implementation.
+This taxonomy clarifies that while prior implementations suffer from multiple bug types, our contribution specifically addresses the batch-specific synchronization challenge that has no analog in single-sequence speculative decoding.
+Appendix B Proof of Synchronization Invariants
+Theorem B.1 (EqSpec: Preservation of Rectangular Alignment).
+Given a batch of size $B$, let $p_{i}^{(t)}$ and $c_{i}^{(t)}$ denote the padding length and content length of sequence $i$ at verification step $t$.
+If the alignment invariant $p_{i}^{(t)}+c_{i}^{(t)}=L^{(t)}$ holds for all $i\in[1,B]$ at step $t$, the EQSPEC update rule guarantees it holds at step $t+1$.
+Proof.
+We prove this by induction on the verification step $t$.
+Base Case ($t=0$): At initialization, all sequences are left-padded to the maximum prompt length $L^{(0)}$. Thus, $\forall i,p_{i}^{(0)}+c_{i}^{(0)}=L^{(0)}$. The invariant holds.
+Inductive Step: Assume the invariant holds at step $t$. During verification, sequence $i$ accepts $a_{i}$ new tokens. The new content length becomes:
+$c_{i}^{(t+1)}=c_{i}^{(t)}+a_{i}$ (4)
+The new target batch length is determined by the longest updated sequence:
+$L^{(t+1)}=\max_{j\in[1,B]}(c_{j}^{(t)}+a_{j})$ (5)
+The EQSPEC algorithm calculates the padding offset $\delta_{i}$ as defined in Eq. (2):
+$\delta_{i}=(L^{(t+1)}-L^{(t)})-a_{i}$ (6)
+The new padding length is updated as $p_{i}^{(t+1)}=p_{i}^{(t)}+\delta_{i}$. We verify the invariant at $t+1$ by substituting these terms:
+$p_{i}^{(t+1)}+c_{i}^{(t+1)}$ $=(p_{i}^{(t)}+\delta_{i})+(c_{i}^{(t)}+a_{i})$
+$=p_{i}^{(t)}+(L^{(t+1)}-L^{(t)}-a_{i})+c_{i}^{(t)}+a_{i}$
+$=(p_{i}^{(t)}+c_{i}^{(t)})-L^{(t)}+L^{(t+1)}$
+By the inductive hypothesis, $p_{i}^{(t)}+c_{i}^{(t)}=L^{(t)}$. Substituting this yields:
+$=L^{(t)}-L^{(t)}+L^{(t+1)}$
+$=L^{(t+1)}$
+Thus, $\forall i,p_{i}^{(t+1)}+c_{i}^{(t+1)}=L^{(t+1)}$. The rectangular alignment is preserved. ∎
+Corollary B.2 (Zero Overhead for Single Sequence).
+For batch size $B=1$, no padding adjustment is required.
+Proof.
+When $B=1$, the batch consists of a single sequence with content length $c_{1}^{(t)}$.
+We can set padding $p_{1}^{(t)}=0$ without loss of generality, satisfying the alignment invariant naturally as $L^{(t)}=c_{1}^{(t)}$.
+Given acceptance $a_{1}$, the new length is $L^{(t+1)}=c_{1}^{(t)}+a_{1}$.
+Substituting into Eq.
+(2):
+$\delta_{1}=(L^{(t+1)}-L^{(t)})-a_{1}=(c_{1}^{(t)}+a_{1}-c_{1}^{(t)})-a_{1}=0$ (7)
+Since $\delta_{1}=0$, no unpad-append-repad or KV-cache shifting is required. ∎
+Corollary B.3 (EXSpec: Zero Overhead for Uniform Acceptance).
+If all sequences in a batch accept the same number of tokens $k\geq 1$, no padding realignment is required ($\delta_{i}=0$ for all $i$), regardless of their initial padding or content lengths.
+Proof.
+Let any sequence $i$ have arbitrary padding $p_{i}^{(t)}$ and content $c_{i}^{(t)}$.
+The batch length at step $t$ is defined as $L^{(t)}=\max_{j}(p_{j}^{(t)}+c_{j}^{(t)})$.
+Recall that the invariant $p_{j}^{(t)}+c_{j}^{(t)}=L^{(t)}$ holds for all sequences.
+If every sequence accepts an identical number of tokens $a_{i}=k$, then the new content length is $c_{i}^{(t+1)}=c_{i}^{(t)}+k$. The new batch length becomes:
+$L^{(t+1)}$ $=\max_{j}(p_{j}^{(t)}+c_{j}^{(t+1)})$
+$=\max_{j}(p_{j}^{(t)}+c_{j}^{(t)}+k)$
+$=(\max_{j}(p_{j}^{(t)}+c_{j}^{(t)}))+k$
+$=L^{(t)}+k$
+The realignment offset $\delta_{i}$ is calculated as:
+$\delta_{i}$ $=(L^{(t+1)}-L^{(t)})-a_{i}$
+$=((L^{(t)}+k)-L^{(t)})-k$
+$=k-k$
+$=0$
+Since $\delta_{i}=0$, the existing padding structure is preserved exactly. No memory reallocation or shifting is required, even if the sequences started with different amounts of padding. ∎
+Appendix C BatchVerify Algorithm Details
+Algorithm 3 implements batch verification through a single forward pass over all draft tokens.
+On the first iteration, the input concatenates the full sequences with their draft tokens ($\mathcal{X}\leftarrow S\oplus D$) to build the initial KV-cache; subsequent iterations process only the draft tokens since prior context is cached.
+The target model produces logits for all positions, from which we extract predicted tokens via argmax.
+Mismatch detection is vectorized: we compare predicted tokens against draft tokens and identify the first disagreement position $j$ per sequence.
+The accepted tokens $A[i]$ consist of all draft tokens up to (but not including) the first mismatch, while the bonus token $B[i]$ is sampled from the target model’s distribution at the mismatch position—critically, this must come from the target model, not the draft model, to preserve output equivalence.
+Note that because bonus token sampling occurs after the forward pass completes, the bonus token lacks KV-cache entries and must be included in the next forward pass.
+An alternative approach, previously employed by vLLM’s v0 engine, sidesteps variable acceptance lengths through batch expansion: rather than handling ragged outputs, each sequence is duplicated $K$ times with progressively longer draft prefixes (e.g., $t_{1}],[t_{1},t_{2}],[t_{1},t_{2},t_{3}]$), all verified in one pass, keeping only the longest correct prefix per sequence.
+While this maintains rectangular tensors, it incurs $K\times$ memory and compute overhead—resources wasted when early tokens are rejected.
+This approach was deprecated in vLLM v1 due to GPU memory overflow at scale and incompatibility with CUDA graphs.
+Our BatchVerify instead accepts the ragged output directly and delegates synchronization to the subsequent unpad-append-repad phase (Section 3.1), avoiding the multiplicative overhead of batch expansion.
+Appendix D Latency-Throughput Tradeoff in Online Serving
+To evaluate performance under realistic online serving conditions, we conducted experiments simulating dynamic request arrivals using full multi-turn conversations from SpecBench.
+We shuffled conversation turns to maximize length diversity and inserted new turns into the next available batch, measuring both throughput (tokens/s) and request completion latency (s) at P50/P90/P99 percentiles.
+Table 4 compares EqSpec against EXSpec across batch sizes 1, 2, 4, and 8.
+BS Method Throughput P50 P90 P99
+1 EqSpec 15.20 6.70 12.96 16.53
+1 EXSpec 15.78 6.08 8.32 9.42
+2 EqSpec 26.70 7.96 14.78 19.06
+2 EXSpec 30.54 9.75 114.89 134.03
+4 EqSpec 46.11 9.46 16.23 19.13
+4 EXSpec 52.44 8.01 53.21 70.35
+8 EqSpec 76.03 11.26 19.23 22.50
+8 EXSpec 77.54 9.16 19.13 33.93
+Table 4: Latency-throughput tradeoff under simulated online serving with heterogeneous request lengths.
+Throughput is in tokens/s.
+P50, P90, and P99 denote the 50th, 90th, and 99th percentiles of request completion latency, respectively.
+Bold values indicate significantly worse tail latencies for EXSpec.
+[…]
+Appendix E Continuous Batching Systems: A Different Problem Domain
+Important Caveat This appendix compares against continuous batching systems (vLLM, SGLang) for completeness, but these systems solve a fundamentally different problem than the static batch speculative decoding addressed in this paper.
+Continuous batching uses variable-length packing with paged attention, which sidesteps the ragged tensor problem entirely through different architectural choices.
+The observations below are not claims that these systems are “incorrect” in the same sense as BSP/DSD—rather, they reflect different design tradeoffs, feature scope limitations, and numerical precision characteristics inherent to their architectures.
+Architectural differences.
+Static batching processes a fixed batch of $B$ sequences together, requiring rectangular tensor alignment (Invariant I1) and facing the ragged tensor synchronization challenge when sequences accept different numbers of draft tokens.
+Continuous batching systems (e.g., vLLM, SGLang) process requests using variable-length packing, sidestepping I1, but still require scatter-gather across requests and rollback both position-ID and KV-cache for rejected tokens—a one-dimensional analog of alignment.
+Critically, I2 remains a prerequisite for algorithmic correctness regardless of batching strategy; note that neither vLLM nor SGLang currently supports continuous batching speculative decoding with external draft models.
+Systems evaluated.
+For informational comparison, we evaluated: vLLM (Kwon et al., 2023) (which subsumes TETRIS (Wu et al., 2025) and TurboSpec (Liu et al., 2025a) as vLLM forks) and SGLang-EAGLE (Zheng et al., 2024; Li et al., 2024a, 2025).
+We used the vLLM v0 engine because v1 deprecates speculative decoding.
+SGLang is compatible only with the EAGLE family; we compare Vicuna-7B/EAGLE2 and Qwen3-8B/EAGLE3 (no available weights for GLM-4).
+We further tested SGLang-EAGLE-Deterministic (SGLang Team, 2025), which enables deterministic execution to reduce numerical variance.
+Algorithm 3 BatchVerify: Single Forward Pass
+0:  Target model $\mathcal{M}_{t}$, Sequences $\mathcal{S}$, Draft tokens $D$, KV cache
+0:  Accepted tokens $A$, Bonus tokens $B$
+1:  if first iteration then
+2:  $\mathcal{X}\leftarrow\mathcal{S}\oplus D$
+3:  else
+4:  $\mathcal{X}\leftarrow D$
+5:  end if
+6:  $\textit{logits},\textit{KVCache}\leftarrow\mathcal{M}_{t}(\mathcal{X},\textit{KVCache})$
+7:  $\textit{pred\_tokens}\leftarrow\arg\max(\textit{logits, dim=vocab})$
+8:  $\triangleright$ Vectorized first mismatch detection
+9:  $\textit{matches}\leftarrow(\textit{pred\_tokens}=D)$
+10:  $J\leftarrow\arg\max(\neg\textit{matches, dim=seq})$
+11:  $\triangleright$ Ragged shape acceptance, no vectorization
+12:  for each sequence $i$ in batch do
+13:  $A[i]\leftarrow D[i][:j]$
+14:  $\triangleright$ Get bonus token from first mismatch
+15:  $\textit{bonus\_logit}\leftarrow\textit{logits}[i,|\mathcal{S}[i]|+j]$
+16:  $B[i]\leftarrow\arg\max(\textit{bonus\_logit})$
+17:  end for
+18:  return $A,B,KVCache$
+Throughput observations.
+[…]
+Method Vicuna Qwen3 GLM4
+E P E P E P
+vLLM + Spec 96.9 98.0 65.6 78.5 72.7 84.5
+SGLang + EAGLE 69.8 79.5 47.7 65.4
+SGLang + EAGLE + Det 85.0 90.0 50.6 69.5
+Table 5: Correctness of continuous batching systems. Exact match (E) and partial match (P) scores compared against each system’s own non-speculative mode.
+Accuracy observations.
+Table 5 shows accuracy metrics for vLLM and SGLang.
+Both systems are accurate on Vicuna but degrade on Qwen3, suggesting model-specific sensitivities to position encoding or numerical precision.
+SGLang-EAGLE-Deterministic (SGLang Team, 2025) helps disambiguate the cause: the improvement from 69.8% to 85.0% on Vicuna confirms that most divergences stem from floating-point non-determinism (He and Lab, 2025) rather than algorithmic bugs.
+Concurrent work by Liu et al.
+(2025b) reaches similar conclusions using output length correlation, though our exact and partial match metrics provide finer-grained verification by distinguishing complete sequence match from partial divergence.
+Note that SGLang only supports EAGLE-family drafters, not external draft models.
+Comparison considerations.
+Direct comparison with these systems is complicated by fundamental architectural differences.
+These frameworks incorporate orthogonal optimizations (CUDA graphs, paged attention, chunked prefilling, continuous batching) independent of speculation.
+Moreover, continuous batching dynamically varies effective batch size based on request load, exposing only a “maximum batch size” parameter rather than a fixed batch size, making controlled comparisons infeasible.
+The deprecation of batch expansion also means neither system currently supports speculative decoding with external draft models at scale.
+Architectural implications.
+The contrast between batch expansion and our approach illustrates two strategies for handling ragged tensors: resource duplication versus explicit synchronization.
+Batch expansion maintains rectangular tensors by duplicating work, trading $K\times$ memory for implementation simplicity.
+Our unpad-append-repad approach instead accepts raggedness and pays synchronization costs only when sequences diverge.
+Integrating this synchronization-based approach with continuous batching and paged attention remains an open problem for production deployment.
+Figure 6: Speculative decoding lags non-speculative baselines, with larger batches further degrading throughput.
+Appendix F Discussion
+Batch parallelism and per-sequence speculative decoding represent two orthogonal acceleration strategies that should, in principle, multiply together: batching exploits GPU parallelism across sequences while speculation reduces sequential token dependencies within each sequence.
+However, when combining these techniques in practice, the ragged tensor problem emerges as a fundamental obstacle that breaks this multiplicative relationship.
+Different sequences accept different numbers of draft tokens, desynchronizing position IDs, attention masks, and KV-cache state across the batch.
+A correctly implemented batch speculative decoding, as we provide through EqSpec and EXSpec, preserves the multiplicative gains but incurs an inherent realignment cost that consumes up to 40% of computation.
+Critically, batch speculation amplifies per-sequence gains; it cannot create them.
+If a draft-target pair shows no speedup at batch size 1 due to low acceptance rates or high draft overhead, batching cannot recover what does not exist at the single-sequence level.
+Our work establishes the synchronization invariants required for algorithmic correctness, quantifies their irreducible costs, and demonstrates through cross-batch scheduling that intelligent system design can mitigate, though not eliminate, the overhead of maintaining output equivalence at scale.
+Appendix G Reproducibility Statement
 We provide our complete implementation in the Supplementary Material, including all hyperparameters, experimental configurations, and model specifications detailed in Section 4.
-Our correctness verification framework using exact and partial match metrics enables deterministic validation of both our results and future implementations.
+Our output equivalence verification framework using exact and partial match metrics enables deterministic validation of both our results and future implementations.
 All experiments use publicly available models and datasets for reproducibility.
@@
 We tested both standard SGLang inference and SGLang with deterministic mode enabled (SGLang Team, 2025) to isolate floating-point non-determinism from algorithmic correctness issues.
-Disclose on LLM usage
+Appendix H Disclose on LLM Usage
 We used Large Language Models as assistive tools in preparing this manuscript: GPT-5 and Claude Opus 4.1 for polishing writing (grammar correction and sentence restructuring), generating conceptual diagram illustrations, and writing unit tests; NVIDIA/Parakeet-TDT-0.6B-v3 for voice input transcription.
@@
 We take full responsibility for all content in this paper.
-Ethics Statement
-This work focuses on improving the efficiency of LLM inference without altering model outputs or introducing biases.
-Our correctness-preserving approach ensures that acceleration techniques do not compromise model safety or reliability.
-Our open-source release enables reproducible research and transparent evaluation of batch speculative decoding techniques.